### v1

# A generalized Schmidt’s subspace theorem for algebraic points of bounded degree

Abstract.
Schmidt’s subspace theorem in terms of Seshadri constants for closed subschemes in subgeneral position has been already developed sharply. We derive our theorem for numerically equivalent ample divisors by dint of the above theory step by step, providing a joint generalization of Schmidt’s theorem for algebraic points of bounded degree.
Key words and phrases:
Schmidt’s subspace theorem, Roth’s theorem, Diophantine approximation, Divisors, Seshadri constants, algebraic point
1. Introduction
1.1. Glorious works for Schmidt’s subspace theorem
Schmidt’s subspace theorem is a fundamental component in the realm of higher-dimensional Diophantine approximation. Over the years, it has been extended and generalized by numerous researchers. Notable contributions have been made by Evertse and Ferretti (as seen in [EF08]), Corvaja and Zannier (as seen in [CZ04a]), and Ru and Vojta [RV20]. A recent version of this theorem, which involves closed subschemes in general position and their Seshadri constants relative to a given ample divisor $A$, has been established. This version builds upon the work of Ru and Wang [RW22] and Vojta [Voj20], who have explored the use of beta constants as an alternative to Seshadri constants. This particular formulation was proven in the recent paper [HL21].
Theorem 1.1 ([HL21, Theorem 1.3]).
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$. Let $S$ be a finite set of places of $k$. For each $v\in S$, let $Y_{0,v},\ldots,Y_{n,v}$ be closed subschemes of $X$ in general position, defined over $k$. Let $A$ be an ample Cartier divisor on $X$. Then there exists a proper Zariski-closed subset $Z\subset X$ such that for all points $P\in X(k)\setminus Z$ and $\epsilon>0$,
$\sum_{v\in S}\sum_{i=0}^{n}\epsilon_{Y_{i,v}}(A)\lambda_{Y_{i,v},v}(P)<(n+1+\epsilon)h_{A}(P).$
More recently, in [Qua22a], Quang introduced the concept of a distributive constant $\Delta$ for a family of divisors, and later in the [Qua22b], Quang also expanded this notable notion to encompass closed subschemes:
Definition 1.2.
For a family of closed subschemes(similar for divisors) $\mathcal{Y}=\{Y_{1},\ldots,Y_{q}\}$ defined on a projective variety $X$, the distributive constant is given by
$\delta_{\mathcal{Y}}=\max_{J\subset\{1,\ldots,q\}}\max\left\{1,\frac{\#J}{\operatorname{codim}\cap_{j\in J}\operatorname{Supp}Y_{j}}\right\},$
where we set $\operatorname{codim}\varnothing=\infty$.
Theorem 1.3 (Quang [Qua22b]).
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $Y_{1,v},\ldots,Y_{q,v}$ be closed subschemes of $X$ defined over $k$. Let $\Delta=\max_{v\in S}\delta_{\mathcal{Y}_{v}}$ be the maximum of the distributive constants of $\mathcal{Y}_{v}=\{Y_{1,v},\ldots,Y_{q,v}\}$. Let $A$ be an ample Cartier divisor on $X$. Then there exists a proper Zariski closed set $Z$ of $X$ such that
$\sum_{v\in S}\sum_{i=1}^{q}\epsilon_{Y_{i,v}}(A)\lambda_{Y_{i,v},v}(P)<\left(\Delta(n+1)+\epsilon\right)h_{A}(P)$
for all points $P\in X(k)\setminus Z$ and $\epsilon>0$.
As discussed in [Qua22a],
$\Delta\leq m-n+1$
if the $Y_{i,v}$ are taken to be closed subschemes in $m$-subgeneral position.
$\Delta\leq\frac{m-n}{\kappa}+1$
if the $Y_{i,v}$ are taken to be closed subschemes in $m$-subgeneral position with index $\kappa$.
Therefore, Theorem 1.3 not only encompasses but also generalizes the prior results concerning closed subschemes in $m$-subgeneral position and its variations.
In the latest paper [HL25], the authors extend Theorem 1.3 to the following Schmidt-Nochka-type Theorem for weighted sums involving arbitrary closed subschemes.
Theorem 1.4 ([HL25], Theorem 1.2).
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $Y_{1,v},\ldots,Y_{q,v}$ be closed subschemes of $X$ defined over $k$, and let $c_{1,v},\ldots,c_{q,v}$ be nonnegative real numbers. For a closed subset $W\subset X$ and $v\in S$, let
$\alpha_{v}(W)=\sum_{\begin{subarray}{c}i\\ W\subset\operatorname{Supp}Y_{i,v}\end{subarray}}c_{i,v}.$
Let $A$ be an ample Cartier divisor on $X$, and $\epsilon>0$. Then there exists a proper Zariski closed subset $Z$ of $X$ such that
$\sum_{v\in S}\sum_{i=1}^{q}c_{i,v}\epsilon_{Y_{i,v}}(A)\lambda_{Y_{i,v},v}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{A}(P)$
for all points $P\in X(k)\setminus Z$.
The application of weights in the inequality described above is entirely innovative.
In comparison with Theorem 1.3, under the assumption $c_{i,v}=1$ for all $i$ and $v$, then
$\delta_{\mathcal{Y}_{v}}=\max\left\{1,\max_{\emptyset\subsetneq W\subsetneq X}\frac{\alpha_{v}(W)}{\operatorname{codim}W}\right\}.$
This conclusion stems directly from the definitions and a straightforward observation (which we will elaborate on in our proof) that when taking the maximum over $W$, it suffices to consider the cases where $W$ is the intersection of a subset of closed subschemes of $\mathcal{Y}_{v}$.
Therefore, as stated in [HL25], Theorem 1.4 can be regarded as a weighted extension of Theorem 1.3 for algebraic points. Specifically, in certain cases, Theorem 1.4 improves upon the coefficient of the height compared to Quang’s theorem, primarily due to the presence of the maximum with 1 in Quang’s definition of $\delta_{\mathcal{Y}}$. And it significantly reduces the discussion requirements for the condition of position. Moreover, this theorem represents the most powerful generalization available to date.
1.2. Constructive results in Schmidt’s theorem for algebraic points of bounded degree
Previous work studying Schmidt’s theorem(for numerically equivalent ample divisors) for algebraic points of bounded degree proved the following theorem,
Theorem 1.5 ([Lev14], Theorem 1.7).
Let $X$ be a projective subvariety of $\mathbb{P}^{N}$ of dimension $n\geq 1$ defined over a number field $k$. Let $S$ be a finite set of places of $k$. For $v\in S$, let $H_{0,v},\ldots,H_{n,v}\subset\mathbb{P}^{N}$ be hypersurfaces over $k$ in n-subgeneral position(general position). Let $\delta\geq 1$ be an integer and $\epsilon>0$. Then the inequality
(1) $\sum_{v\in S}\sum_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\sum_{i=0}^{n}\frac{\lambda_{H_{i,v},w}(P)}{\deg H_{i,v}}<\left((\delta n)^{2}\left(\frac{\delta n-1}{2\delta n-3}\right)+\epsilon\right)h(P)$
holds for all but finitely many points $P\in X(\overline{k})\setminus\cup_{v\in S}\cup_{i=0}^{n}H_{i,v}$ satisfying $[k(P):k]\leq\delta$.
Set $X=\mathbb{P}^{n}=\mathbb{P}^{N}$, while the $H_{i,v}\subset X$ are hyperplanes, Schlickewei proposed that there exists a constant $c(\delta,n)$, which depends solely on $\delta$ and $n$, and a finite union of hyperplanes $Z\subset\mathbb{P}^{n}$ such that the left-hand side of inequality (1) is bounded by $(c(\delta,n)+\epsilon)h(P)$ for all points $P\in\mathbb{P}^{n}(\overline{k})\setminus Z$ satisfying $[k(P):k]\leq\delta$. Consequently, Theorem 1.5 confirms Schlickewei’s conjecture as a special case.
Combined with the inequality we will prove more detailed, if $\delta m\geq 2$, we have a more general theorem(Theorem 1.6) proved in [Lev14]
$C(\delta,1,m,n)\leq\max_{1\leq j\leq\delta n}A(\delta m,j)\leq\max_{1\leq j\leq\delta n}\frac{\delta m(\delta m-1)(j+1)}{\delta m+j-2}$
$\leq$ $\frac{\delta m(\delta m-1)(\delta n+1)}{\delta m+\delta n-2}$
Explicitly, this provides an adaptation of Schmidt’s theorem for algebraic points.
Theorem 1.6 ([Lev14], Theorem 5.2).
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$. Let $S$ be a finite set of places of $k$. For each $v\in S$, let $D_{0,v},\ldots,D_{m,v}$ be effective ample Cartier divisors on $X$ in $m$-subgeneral position, defined over $k$. Suppose that there exists a Cartier divisor $A$ on $X$ defined over $k$, and positive integers $d_{i,v}$ such that $D_{i,v}\equiv d_{i,v}A$ for all $i$ and for all $v\in S$. Let $\delta$ be a positive integer with $\delta m\geq 2$ and $\epsilon>0$. Then the inequality
$\sum_{v\in S}\sum_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\sum_{i=0}^{m}\frac{\lambda_{D_{i,v},w}(P)}{d_{i,v}}<\left(\frac{\delta m(\delta m-1)(\delta n+1)}{\delta m+\delta n-2}+\epsilon\right)h_{A}(P)+O(1)$
holds for all points $P\in X(\overline{k})\setminus\cup_{i,v}\operatorname{Supp}D_{i,v}$ satisfying $[k(P):k]\leq\delta$.
If $m=n\geq 2$, then using $A(\delta n,\delta n)\leq\delta n+1$ gives the following slightly better bound(which is essentially quadratic in $\delta n$) compared with the previous conclusions:
$C(\delta,1,n,n)\leq\max_{1\leq j\leq\delta n}A(\delta n,j)\leq(\delta n)^{2}\frac{\delta n-1}{2\delta n-3}.$
This gives Theorem 1.5, as a special case.
Comprehensively, inspired by the newest work in Schmidt’s subspace theorem, the following theorem is our main result:
Theorem 1.7.
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $D_{1,v},\ldots,D_{q,v}(q=m+1)$ be effective ample Cartier divisors of $X$ defined over $k$, let $c_{i,v}=1(1\leq i\leq q)$ and set Seshadri constants $\epsilon_{D_{i,v}}(A)$ construct positive integers $d_{i,v}$ such that $D_{i,v}\equiv d_{i,v}A$ for all $i$ and for all $v\in S$. For a closed subset $W\subset X$ and $v\in S$, let
$\alpha_{v}(W)=\sum_{\begin{subarray}{c}i\\ W\subset\operatorname{Supp}D_{i,v}\end{subarray}}c_{i,v}.$
Let $A$ be an ample Cartier divisor on $X$, and $\epsilon>0$. Then
$\sum_{v\in S}\sum_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\sum_{i=1}^{q}\epsilon_{D_{i,v}}(A)\lambda_{D_{i,v},w}(P)<\left(a_{\delta}+\epsilon\right)h_{A}(P)+O(1)$
where
$a_{\delta}=\left[(n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}\right]_{w.r.t-\delta}$
holds for all points $P\in X(\overline{k})\setminus\cup_{i,v}\operatorname{Supp}D_{i,v}$ satisfying $[k(P):k]\leq\delta$, $\delta$ is a positive integer.
$[*]_{w.r.t-\delta}$ here is a consequence of an operation by dint of the method in our proof section. In view of the particularity of various situations, we will systematically illustrate and list the results in the following remark and uniformly use $m$-subgeneral position and $c_{i,v}=1$ as the main research object to get trivial solution, which is equivalent to a sharp integration of Levin’s results in [Lev14].
Remark 1.8.
In the case of $m$-subgeneral position and the same assumption as above, let $X\subset\mathbb{P}^{N}(k)$ be a projective variety of dimension $n$ defined over a number field $k$, if $D$ is a hypersurface in $\mathbb{P}^{N}(k)$ defined by a homogeneous polynomial $f\in k[x_{0},\ldots,x_{N}]$ of degree $d$, for $v\in M_{k}$, we let $|f|_{v}$ denote the maximum of the absolute values of the coefficients of $f$ with respect to $v$. We define $\|f\|_{v}$ similarly. then a Weil function for $D$ is given by
$\lambda_{D,v}(P)=\log\frac{\|f\|_{v}\max_{i}\|x_{i}\|_{v}^{d}}{\|f(P)\|_{v}}.$
so for $\mathcal{D}_{v}=\{D_{1,v},\ldots,D_{q,v}\}(q=m+1)$ be hypersurfaces, we have,
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{I=1}^{q}\left(\frac{\|f_{I}\|_{v,w}\max_{0\leq i\leq N}\|x_{I_{i}}\|^{d_{I}}_{v,w}}{\|f_{I}(P)\|_{v,w}}\right)^{\epsilon_{D_{I,v}}(A)}<\left(a_{\delta}+\epsilon\right)h_{A}(P)+O(1)$
holds for all points $P\in X(\bar{k})\setminus\cup_{I,v}\operatorname{Supp}D_{I,v}$ satisfying $[k(P):k]\leq\delta$.
By our assumption and discussion in proof section and due to $c_{I,v}=1$ for all $I$(that’s the point we focus on and the equation above is equally applicable to numerically equivalent ample divisors) we modify $a_{\delta}$ into $a_{\Delta,\delta}$, where $a_{\Delta,\delta}$ here is defined by
$\left[(n+1)\max_{v\in S}\delta_{\mathcal{D}_{v}}(i.e-\Delta)\right]_{w.r.t-\delta}$
we get the following main results:
(i)For $m>2n$,
$a_{\Delta,\delta}=n({\delta^{2}}m-{\delta^{2}}n)+{\delta}m+1$
(ii)For $m\leq 2n$, set $[*]_{Int}$ as a rounding symbol, taking the nearest integer to *(or just identity), then for $j(1\leq j\leq\delta n)$ as above, $j=\left[\frac{\delta m}{2}\right]_{Int}$ in this case, so we have
$a_{\Delta,\delta}=-j^{2}+\delta mj+{\delta}m+1$
This improved an earlier result in [Lev14] which had a factor of
$\frac{\delta m(\delta m-1)(\delta n+1)}{\delta m+\delta n-2}+\epsilon$
on the right-hand side. Furthermore, for more general case $m$-subgeneral position with index $\kappa$, we have
(i)For $m>\lceil 2n+\frac{1-\kappa}{\delta}\rceil$,
$a_{\Delta,\delta}=\frac{n}{\kappa}\left({\delta^{2}}m-{\delta^{2}}n+\delta(\kappa-1)\right)+\frac{{\delta}m}{\kappa}+1$
(ii)For $m\leq\lfloor 2n+\frac{1-\kappa}{\delta}\rfloor$, as above, $j=\left[\frac{\delta m+(\kappa-1)}{2}\right]_{Int}$ in this case, then
$a_{\Delta,\delta}=-\frac{j^{2}}{\kappa}+(\frac{\delta m}{\kappa}-\frac{1}{\kappa}+1)j+\frac{{\delta}m}{\kappa}+1$
The structure of this paper is as follows. In Section 2, we recall some basic definitions and introduce some notions. In Section 3, we will divide the problem into three parts, initially provide a rigorous proof of Schmidt’s subspace theorem for effective Cartier divisors based on the idea in [HL25] and extent it to numerically equivalent ample divisors, and then handle the case of algebraic point in terms of $m$-subgeneral position with $c_{i,v}=1$ to facilitate more improvements on discussions in [Lev14].
2. Definitions and preliminaries
Throughout this paper, we denote $k$ as a number field and $\mathcal{O}_{k}$ as the ring of integers of $k$. We define $M_{k}$ as a canonical set of places (or absolute values) of $k$, which includes: (i) One place for each prime ideal $\mathfrak{p}$ of $\mathcal{O}_{k}$, (ii) One place for each real embedding $\sigma:k\to\mathbb{R}$, (iii) One place for each pair of conjugate embeddings $\sigma,\overline{\sigma}:k\to\mathbb{C}$.
Given a finite set $S$ of places of $k$ that contains all the archimedean places, we denote $\mathcal{O}_{k,S}$ as the ring of $S$-integers of $k$, $\mathcal{O}_{k,S}^{*}$ as the group of $S$-units of $k$. For a place $v$ of $k$ and a place $w$ of a field extension $L$ of $k$, we say that $w$ lies above $v$(denoted $w|v$) if $w$ and $v$ induce the same topology on $k$. We normalize the absolute values as follows: $|p|_{v}=\frac{1}{p}$ if $v$ corresponds to a prime ideal $\mathfrak{p}$ lying above a rational prime $p$, $|x|_{v}=|\sigma(x)|$ if $v$ corresponds to an embedding $\sigma$. For each $v\in M_{k}$, we let $k_{v}$ denote the completion of $k$ with respect to $v$. We set
$\|x\|_{v}=|x|_{v}^{[k_{v}:\mathbb{Q}_{v}]/[k:\mathbb{Q}]}.$
A fundamental equation is the product formula
$\prod_{v\in M_{k}}\|x\|_{v}=1,$
which holds for all $x\in k^{*}$.
We start by providing a concise overview of the essential properties of local and global height functions, primarily following the reference [Sil87] and [HL25]. Subsequently, we will introduce the concept of subgeneral position and Seshadri constants, which are utilized in this paper.
For more refined versions and more details about Local(Global) height functions can be found in [HL25], we will only provide a brief explanation here.
2.1. Local height functions
Let $Y$ be a closed subscheme of a projective variety $X$, both defined over a number field $k$. For any place $v$ of $k$, one can associate a local height function (or Weil function) $\lambda_{Y,v}:X(k)\setminus Y\to\mathbb{R}$, up to a bounded error $O(1)$. This function measures the $v$-adic distance of a point in $X(k)\setminus Y$ to the closed subscheme $Y$, taking large values when the point is close to $Y$. It satisfy the following properties: if $Y$ and $Z$ are two closed subschemes of $X$, defined over $k$, and $v$ is a place of $k$, then up to $O(1)$,
$\lambda_{Y\cap Z,v}=\min\{\lambda_{Y,v},\lambda_{Z,v}\},\lambda_{Y+Z,v}=\lambda_{Y,v}+\lambda_{Z,v},\lambda_{Y,v}\leq\lambda_{Z,v},\text{ if }Y\subset Z.$
In particular, $\lambda_{Y,v}$ is bounded from below for all $P\in X(k)\setminus Y$. If $\phi:W\to X$ is a morphism of projective varieties with $\phi(W)\not\subset Y$, then up to $O(1)$,
$\lambda_{Y,v}(\phi(P))=\lambda_{\phi^{*}Y,v}(P),$
$\quad\forall P\in W(k)\setminus\phi^{*}Y.$
The operation $Y\cap Z$, $Y+Z$, $Y\subset Z$, and $\phi^{*}Y$ are defined in terms of the associated ideal sheaves. Specifically, it is a subtle point that if $Y$ is associated with the ideal sheaf $\mathcal{I}_{Y}$, then $\phi^{*}Y$ represents the closed subscheme corresponding to the inverse image ideal sheaf $\phi^{-1}\mathcal{I}_{Y}\cdot\mathcal{O}_{W}$. If $Y=D$ is an effective (Cartier) divisor (which we often identify with the associated closed subscheme), these height functions coincide with the usual height functions associated to divisors.
It is worth highlighting that by setting $Y=Z$, we obtain the following equality
$\lambda_{2Y,v}=\lambda_{Y+Y,v}=\lambda_{Y,v}+\lambda_{Y,v}=2\lambda_{Y,v},$
where $2Y$ denotes the closed subscheme corresponding to the ideal sheaf $\mathcal{I}_{Y}^{2}$. Furthermore, it is generally true that if $Y_{1},\ldots,Y_{m}$ are closed subschemes of $X$ and $c_{1},\ldots,c_{m}$ are positive integers, we have
(2) $\lambda_{c_{1}Y_{1}\cap\ldots\cap c_{m}Y_{m},v}$ $=\min\{c_{1}\lambda_{Y_{1},v},\ldots,c_{m}\lambda_{Y_{m},v}\},$
(3) $\lambda_{c_{1}Y_{1}+\cdots+c_{m}Y_{m},v}$ $=c_{1}\lambda_{Y_{1},v}+\cdots+c_{m}\lambda_{Y_{m},v}.$
2.2. Global height functions
As our research focus, global height functions of Cartier divisors, especially ample Cartier divisors, play a crucial role. Same definition for $Y$ and $X$, then a global height function $h_{Y}:X(k)\setminus Y\to\mathbb{R}$, can be associated as follows,
$h_{Y}(P)=\sum_{v\in M_{k}}\lambda_{Y,v}(P).$
Where $M_{k}$ denote the set of all places of $k$. For $v\in M_{k}$, the definition of $\lambda_{Y,v}$ can be chosen such that the collection $(\lambda_{Y,v})_{v\in M_{k}}$ is well-defined up to an $M_{k}$-constant($M_{k}$-constant is a sequence of real numbers $(c_{v})_{v\in M_{K}}$ where $c_{v}=0$ for all but finitely many $v$). As a result, $h_{Y}$ is well-defined up to $O(1)$.
These functions possess properties similar to those described earlier for local height functions, with one notable exception: the property above is modified to $h_{Y\cap Z}\leq\min\{h_{Y},h_{Z}\}+O(1)$. All height function equalities and inequalities implicitly account for bounded ambiguity, even if not explicitly stated or if references to $M_{k}$-constants or $O(1)$ terms are omitted. This is a standard convention in the field.
Specifically, for an ample Cartier divisor $A$, the global height function $h_{A}$ quantifies the arithmetic complexity of a point $P$. Additionally, for any arbitrary closed subscheme $Y$, the height function $h_{A}$ dominates $h_{Y}$, as the domain of $h_{A}$ can be extended to encompass all of $X(k)$ and there exists a constant $c$ such that
(4) $h_{Y}(P)\leq ch_{A}(P)$
for all $P\in X(k)\setminus Y$. The reference [Sil87, Proposition 4.2] explains that the intuition behind this result is as follows: if a point $P$ is $v$-adically close to $Y$(though not on $Y$), then after a projective embedding using a very ample multiple of the divisor $A$, the coordinates of $P$ become $v$-adically complex.
Another important property of the local height functions ($\lambda_{Y,v})_{v\in M_{k}}$ is that they are bounded from below by an $M_{k}$-constant. This fact immediately implies that the inequality (4) can be formulated for each individual $\lambda_{Y,v}$, i.e.,
(5) $\lambda_{Y,v}(P)\leq ch_{A}(P).$
We will all use the above properties in our following proof.
We further define the vector space $L(D)=\{f\in\overline{k}(X)\mid\text{div}(f)\geq-D\}$, where $\text{div}(f)$ denotes the divisor associated with $f$. The dimension of this vector space over $\overline{k}$ is denoted by $l(D)=\dim_{\overline{k}}L(D)$. We use the symbol $\sim$ to indicate linear equivalence of divisors and $\equiv$ to denote numerical equivalence of divisors.
2.3. Subgeneral position and Seshadri constants
We review the definitions of position and of Seshadri constants, both in the context of closed subschemes (with analogous definitions applying to divisors).
Definition 2.1.
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$ and let $Y_{1},\ldots,Y_{q}$ be $q$ closed subschemes of $X.$ Suppose $m$ and $\kappa$ are positive integers satisfying $m\geq\dim X\geq\kappa.$
(a). The closed subschemes $\{Y_{1},\ldots,Y_{q}\}$ are taken to be in general position in $X$ if for any subset $I\subset\{1,\ldots,q\}$ with $\sharp I\leq\dim X+1,$
$\operatorname{codim}\left(\bigcap_{i\in I}Y_{i}\cap X\right)\geq\sharp I.$
(b). The closed subschemes $\{Y_{1},\ldots,Y_{q}\}$ are taken to be in $m$-subgeneral position in $X$ if for any subset $I\subset\{1,\ldots,q\}$ with $\sharp I\leq m+1,$
$\dim\left(\bigcap_{i\in I}Y_{i}\cap X\right)\leq m-\sharp I.$
(c). The closed subschemes $\{Y_{1},\ldots,Y_{q}\}$are in $m-$subgeneral position in $X$ with index $\kappa$ if $\{Y_{1},\ldots,Y_{q}\}$ are in $m$-subgeneral position and for any subset $I\subset\{1,\ldots,q\}$ with $\sharp I\leq\kappa,$
$\operatorname{codim}\left(\bigcap_{i\in I}Y_{i}\cap X\right)\geq\sharp I.$
We now revisit the concept of a Seshadri constant for a closed subscheme relative to a nef Cartier divisor (for additional details, see [HL21, Section 2]).
2.4. Seshadri constants
Let $X$ be a projective variety and $Y$ be a closed subscheme of $X,$ corresponding to a coherent sheaf of ideals $\mathcal{I}_{Y}.$ Consider the sheaf of graded algebras $\mathcal{S}=\oplus_{d\geq 0}\mathcal{I}_{Y}^{d}$, where $\mathcal{I}_{Y}^{d}$ is the $d$-th power of $\mathcal{I}_{Y},$ and $\mathcal{I}_{Y}^{0}=\mathcal{O}_{X}$ by convention. Then $\tilde{X}:=Proj\mathcal{S}$ is called the blowing-up of $X$ with respect to $\mathcal{I}_{Y},$ in other words, the blowing-up of $X$ along $Y.$
Let $\pi:\tilde{X}\to X$ be the blow-up of $X$ along $Y$. The inverse image ideal sheaf ~$\tilde{\mathcal{I}}_{Y}=\pi^{-1}\mathcal{I}_{Y}\cdot\mathcal{O}_{\tilde{X}}$ is an invertible sheaf on $\tilde{X}$. Let $E$ be an effective Cartier divisor on $\tilde{X}$ such that the associated invertible sheaf is the dual of $\pi^{-1}\mathcal{I}_{Y}\cdot\mathcal{O}_{\tilde{X}}$.
Definition 2.2.
Let $Y$ be a closed subscheme of a projective variety $X$ and let $\pi:\tilde{X}\to X$ be the blowing-up of $X$ along $Y$. Let $A$ be a nef Cartier divisor on $X$. The Seshadri constant $\epsilon_{Y}(A)$ of $Y$ with respect to $A$ is the real number defined by
[…]
where $E$ is an effective Cartier divisor on $\tilde{X}$ whose associated invertible sheaf is the dual of $\pi^{-1}\mathcal{I}_{Y}\cdot\mathcal{O}_{\tilde{X}}$.
3. Proof of Theorem 1.7 and Remark 1.8
In this section we disassemble our proof precisely into the following implications(by making use of methods of Levin[Lev14]):
  1. (a)
    Schmidt’s theorem for numerically equivalent ample divisors
    $\Downarrow$
  2. (b)
    (a) + controlling the dimension of the exceptional set $Z$
    $\Downarrow$
  3. (c)
    (b) for algebraic points of bounded degree
After proving the theorem of type (a), we can derive the results stated in (b) and (c). Specifically, the implication from (a) $\Rightarrow$ (b) is achieved through repeated applications of (a) to exceptional sets. Meanwhile, the implication from (b) $\Rightarrow$ (c) is established by reducing Diophantine approximation problems for algebraic points of bounded degree on $X$ to those for rational points on symmetric powers of $X$.
First we introduce some parameters that we need, set Seshadri constants $\epsilon_{D_{i,v}}(A)$ construct $d_{i,v}$ one-to-one in our following proof. Continuing to use the symbol system in [Lev14].
Given the setup, we are considering a projective variety $X$ of dimension $n$ defined over a number field $k$, with a finite set of places $S$ of $k$. Let $m\geq n$ be positive integers. For each place $v\in S$, we have $m+1$ effective Cartier divisors $D_{0,v},D_{1,v},\ldots,D_{m,v}$ on $X$ that are in $m$-subgeneral position, i.e., $\cap_{i=0}^{m}\operatorname{Supp}D_{i,v}=\varnothing$. We define a divisor $D$ as the sum of all these divisors over all places in $S$, i.e., $D=\sum_{v\in S}\sum_{i=0}^{m}D_{i,v}$. Additionally, we are given an ample Cartier divisor $A$ on $X$ and positive integers $d_{i,v}$ such that each divisor $D_{i,v}$ is numerically equivalent to $d_{i,v}A$, i.e., $D_{i,v}\equiv d_{i,v}A$ for all $i$ and $v\in S$. For $c\in\mathbb{R}$ and $\delta\in\mathbb{N}$, define the geometric exceptional set $Z=Z\left(c,\delta,m,n,k,S,X,A,\{D_{i,v}\}\right)$ to be the smallest Zariski-closed subset of $X$ such that the set
(6) $\left\{P\in X(\overline{k})\setminus\operatorname{Supp}D\mid\sum_{v\in S}\sum_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\sum_{i=0}^{m}\epsilon_{D_{i,v}}(A)\lambda_{D_{i,v},w}(P)>ch_{A}(P)\right\}\setminus Z$
satisfied for $[k(P):k]\leq\delta$, is finite for all choices of the Weil and height functions.
Now define
$C(\delta,l,m,n)=\sup\{c\mid\dim Z\left(c,\delta,m,n,k,S,X,A,\{D_{i,v}\}\right)\geq l\\ \text{ for some $k$, $S$, $X$, $A$, $\{D_{i,v}\}$ as above}\}.$
Furthermore, it would be beneficial to define and examine specific instances concerning rational points.
$A(m,n)=C(1,n,m,n),B(l,m,n)=C(1,l,m,n).$
In order to prove Theorem 1.7, we require the following inequality due to [HL25](The proof process is consistent, and we will not elaborate further here), which is a generalization of Chebyshev’s inequality (see also work of Jensen [MPF93, p. 245]).
Lemma 3.1.
Let Weil functions family $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}\geq 0$, set $\lambda_{i}=\log\boxdot_{i}$ and $b_{1},\ldots,b_{n},c_{1},\ldots,c_{n}$ be nonnegative real numbers as sets of indicators. Let $i_{0}$ be the smallest one of the index $i$ such that $c_{i}\not=0$(assume that there exists at least one satisfied), then
(7) $\log\prod_{i=1}^{n}\boxdot^{b_{i}}_{i}\geq\log\prod_{i=1}^{n}\boxdot^{\left(\min_{i_{0}\leq j\leq n}\frac{\sum_{i=1}^{j}b_{i}}{\sum_{i=1}^{j}c_{i}}\right)c_{i}}_{i}.$
An immediate corollary of Lemma 3.1 is
Corollary 3.2.
Let Weil functions family $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}\geq 0$, set $\lambda_{i}=\log\boxdot_{i}$ and $b_{1},\ldots,b_{n},c_{1},\ldots,c_{n}$ be nonnegative real numbers as sets of indicators. Assume that $b_{1}\not=0$. Then
$\log\prod_{i=1}^{n}\boxdot^{\left(\max_{1\leq j\leq n}\frac{\sum_{i=1}^{j}c_{i}}{\sum_{i=1}^{j}b_{i}}\right)b_{i}}_{i}\geq\log\prod_{i=1}^{n}\boxdot^{c_{i}}_{i}.$
In order to better elaborate the Remark 1.8, we will use the full representation of Weil function in our following proof.
We now conduct the proof of Theorem 1.7(using the same method in [HL25] to attain the result of effective(Cartier) divisors) and Remark 1.8.
Proof.
Part I: Schmidt’s theorem for numerically equivalent ample divisors.
Let $v\in S$, $c_{i,v}(1\leq i\leq q)$ be nonnegative real numbers. Let $D_{i,v}$ be effective(Cartier) divisors.
(i) Normalization in two step process(In view of the Seshadri constants $\epsilon_{D_{1,v}}(A),\ldots,\epsilon_{D_{q,v}}(A)$ may be irrelevant distinct positive real numbers):
(a) Rationalization: firstly for every $i=1\ldots,q$, let $\tau_{1,v},\ldots,\tau_{q,v}$ be real numbers in the interval $[0,1)$ so that
$(1+\tau_{1,v})\epsilon_{D_{1,v}}(A),\ldots,(1+\tau_{q,v})\epsilon_{D_{q,v}}(A)$
are rational numbers. If $\epsilon_{D_{i,v}}(A)$ is already a rational number then choose $\tau_{i,v}=0$.
(b) Integer: For a sufficiently divisible integer $c_{v}$ (independent of $i$) such that
$c_{v}(1+\tau_{1,v})\epsilon_{D_{1,v}}(A),\ldots,c_{v}(1+\tau_{q,v})\epsilon_{D_{q,v}}(A)$
are all integers.
By means of the multiplication of the effective(Cartier) divisors $D_{i,v}$, we set
$\tilde{D}_{i,v}:=c_{v}(1+\tau_{i,v})\epsilon_{D_{i,v}}(A)D_{i,v},$
we have that
(8) $\epsilon_{\tilde{D}_{i,v}}(A)=\frac{1}{c_{v}(1+\tau_{i,v})}$
for $i=1,\ldots,q$, and therefore
$\frac{1}{c_{v}}\leq 2\epsilon_{\tilde{D}_{i,v}}(A)$
for $i=1,\ldots,q$. Consequently,
$\frac{1}{c_{v}}\leq 2\min_{i}\epsilon_{\tilde{D}_{i,v}}(A).$
For each Seshadri constant and for all $i,i^{\prime}=1,\ldots,q$, we have
$\ |\epsilon_{\tilde{D}_{i,v}}(A)-\epsilon_{\tilde{D}_{i^{\prime},v}}(A)|$
$=$ $\ \left|\frac{1}{c_{v}(1+\tau_{i,v})}-\frac{1}{c_{v}(1+\tau_{i^{\prime},v})}\right|$
$=$ $\ \frac{1}{c_{v}}\left|\frac{1}{(1+\tau_{i,v})}-\frac{1}{(1+\tau_{i^{\prime},v})}\right|$
$=$ $\ \frac{1}{c_{v}}\left|\frac{\tau_{i^{\prime},v}-\tau_{i,v}}{(1+\tau_{i^{\prime},v})(1+\tau_{i,v})}\right|$
$\leq$ $\ \max_{k}\frac{\tau_{k,v}}{c_{v}}$
(9) $\leq$ $\ \max_{k}\tau_{k,v}\cdot 2\min_{l}\epsilon_{\tilde{D}_{l,v}}(A)$
(ii) Intersections of the normalized divisors: set
$\tilde{D}_{I,v}=\bigcap_{i\in I}\tilde{D}_{i,v}.$
Fix a point $P$ outside of the support of the given divisors. Let $\{i_{1,v},\ldots,i_{q,v}\}=\{1,\ldots,q\}$ be such that:
(10) $\log\boxdot_{\tilde{D}_{i_{1,v},v},v}(P)\geq\ldots\geq\cdots\geq\log\boxdot_{\tilde{D}_{i_{q,v},v},v}(P)\geq 0.$
We set $I_{j,v}=\{i_{1,v},\ldots,i_{j,v}\}$. Let $l_{v}$ be the largest index such that
$\cap_{j=1}^{l_{v}}\tilde{D}_{i_{j,v},v}\neq\varnothing.$
With the property (2) and the sequence of inequalities (10) we have
$\log\boxdot_{\tilde{D}_{i_{j,v}v},v}(P)=\min_{j^{\prime}\leq j}\log\boxdot_{\tilde{D}_{i_{j^{\prime},v},v},v}(P)=\log\boxdot_{\tilde{D}_{I_{j,v},v},v}(P).$
In particular, for $j>l_{v}$, $\log\boxdot_{\tilde{D}_{i_{j,v},v},v}(P)$ is bounded (by a constant independent of $P$) and
$\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot^{c_{i,v}\epsilon_{\tilde{D}_{i,v}}(A)}_{\tilde{D}_{i,v},v}(P)\leq\log\prod_{v\in S}\prod_{j=1}^{l_{v}}\boxdot^{c_{i_{j},v}\epsilon_{\tilde{D}_{i_{j,v},v}}(A)}_{\tilde{D}_{I_{j,v},v},v}(P).$
Let $b_{j,v}=\operatorname{codim}\tilde{D}_{I_{j,v},v}$ and set $b_{0,v}=0$. Set $\epsilon_{v}(A)=\max_{i}\epsilon_{\tilde{D}_{i,v}}(A)$. Apply Corollary 3.2 in this condition with
$\lambda_{1}$ $=\log\boxdot^{\epsilon_{v}(A)}_{\tilde{D}_{I_{1,v},v},v}(P)\geq\cdots\geq\lambda_{l_{v}}=\log\boxdot^{\epsilon_{v}(A)}_{\tilde{D}_{I_{l_{v},v},v},v}(P)\geq 0,$
$b_{1}$ $=b_{1,v}-b_{0,v},\ldots,b_{l_{v}}=b_{l_{v},v}-b_{l_{v}-1,v},$
$c_{j}$ $=c_{i_{j},v}.$
Using the telescoping property for the $b_{i}$,
(11) $\log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{\left(\max_{j}\frac{\sum_{j^{\prime}=1}^{j}c_{i_{j^{\prime}},v}}{b_{j,v}}\right)(b_{s,v}-b_{s-1,v})\epsilon_{v}(A)}_{\tilde{D}_{I_{s,v},v},v}(P)\geq\log\prod_{v\in S}\prod_{j=1}^{l_{v}}\boxdot^{c_{i_{j},v}\epsilon_{v}(A)}_{\tilde{D}_{I_{j,v},v},v}(P).$
According to [Laz04, Example 5.4.11] (This is expressed in terms of the s -invariant, which is the reciprocal of the Seshadri constant), the Seshadri constant of an intersection is bounded below by the minimum of the Seshadri constants of the factors of the intersection. Therefore, for any $j\in\{1,\ldots,l_{v}\}$, combined with the (9), we have the inequality
(12) $\epsilon_{\tilde{D}_{I_{j,v},v}}(A)\geq\min\{\epsilon_{\tilde{D}_{i_{1},v}}(A),\ldots,\epsilon_{\tilde{D}_{i_{j},v}}(A)\}\geq\epsilon_{v}(A)-\max_{k}\tau_{k,v}\cdot 2\min_{l}\epsilon_{\tilde{D}_{l,v}}(A),$
Then (11) implies
$\log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{\left(\max_{j}\frac{\sum_{j^{\prime}=1}^{j}c_{i_{j^{\prime}},v}}{b_{j,v}}\right)(b_{s,v}-b_{s-1,v})\left(\epsilon_{\tilde{D}_{I_{s,v},v},v}(A)+\max_{k}\tau_{k,v}\cdot 2\min_{l}\epsilon_{\tilde{D}_{l,v}}(A)\right)}_{\tilde{D}_{I_{s,v},v},v}(P)$
$\geq$ $\log\prod_{v\in S}\prod_{j=1}^{l_{v}}\boxdot^{c_{i_{j},v}\epsilon_{v}(A)}_{\tilde{D}_{I_{j,v},v},v}(P)\geq\log\prod_{v\in S}\prod_{j=1}^{l_{v}}\boxdot^{c_{i_{j},v}\epsilon_{\tilde{D}_{i_{j,v},v}}(A)}_{\tilde{D}_{I_{j,v},v},v}(P).$
If we construct a new list by repeating the divisors $\tilde{D}_{I_{s,v},v}$ exactly $b_{s,v}-b_{s-1,v}$ times (and omitting them if $b_{s,v}-b_{s-1,v}=0$), then the resulting divisors will be in general position. By Theorem 1.1 in the case of effective(Cartier) divisors, there exists a proper Zariski closed subset $Z$ of $X$ such that
$\log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{(b_{s,v}-b_{s-1,v})\epsilon_{\tilde{D}_{I_{s,v},v}}(A)}_{\tilde{D}_{I_{s,v},v},v}(P)<(n+1+\epsilon^{\prime})h_{A}(P)$
for all points $P\in X(k)\setminus Z$. We also find that by (8) and additivity for local heights, for $i=1,\ldots,q$, we have
$\log\boxdot^{\epsilon_{\tilde{D}_{i,v}}(A)}_{\tilde{D}_{i,v},v}(P)=\log\boxdot^{\frac{1}{c_{v}(1+\tau_{i,v})}\cdot c_{v}(1+\tau_{i,v})\epsilon_{D_{i,v}}(A)}_{D_{i,v},v}(P)=\log\boxdot^{\epsilon_{D_{i,v}}(A)}_{D_{i,v},v}(P).$
For a given $\epsilon^{\prime\prime}>0$, by selecting the quantities $\tau_{k,v}$ sufficiently small (with the choice depending only on data involving the $D_{i,v}$, including the finite number of pertaining domination inequalities of the form (5) for the $D_{i,v}$), we get
$\log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{(b_{s,v}-b_{s-1,v})(\max_{k}\tau_{k,v}\cdot 2\min_{l}\epsilon_{\tilde{D}_{l,v}}(A))}_{\tilde{D}_{I_{s,v},v},v}(P)$
$\leq$ $\ \log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{(b_{s,v}-b_{s-1,v})(\max_{k}\tau_{k,v}\cdot 2\epsilon_{\tilde{D}_{i_{s,v},v}}(A))}_{\tilde{D}_{i_{s,v},v},v}(P)$
$\leq$ $\ \log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{(b_{s,v}-b_{s-1,v})(\max_{k}\tau_{k,v}\cdot 2\epsilon_{D_{i_{s,v},v}}(A))}_{D_{i_{s,v},v},v}(P)$
$<$ $\ \epsilon^{\prime\prime}h_{A}(P).$
Taking $W=\operatorname{Supp}\tilde{D}_{I_{j,v},v}\neq\varnothing$, we have
$\frac{\sum_{j^{\prime}=1}^{j}c_{i_{j^{\prime}},v}}{b_{j,v}}\leq\frac{\alpha_{v}(W)}{\operatorname{codim}W}.$
It should be noted that while the indices $i_{1,v},\ldots,i_{l_{v},v}$ for $v\in S$, depended on the point $P$, there are only finitely many possible combinations for these indices. Consequently, only finitely many applications of Theorem 1.1 are required. Taking all of the above into consideration, there exists a proper Zariski closed set $Z$ of $X$ such that
(13) $\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot^{c_{i,v}\epsilon_{D_{i,v}}(A)}_{D_{i,v},v}(P)\leq\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{A}(P)$
for all points $P\in X(k)\setminus Z$, as desired.
Now it shifts to proving linear equivalent divisors:
Let $N$ be a positive integer satisfying $NA$ is very ample and $d_{i,v}$(constructed by $\epsilon_{D_{i,v}}(A)$) divides $N$ for all $i$ and $v\in S$. Define $M=l(NA)-1$, where $l(NA)$ donates the dimension of the space of global sections of $NA$. Let $\phi:X\to\mathbb{P}^{M}$ be the embedding of $X$ in $\mathbb{P}^{M}$ associated to $NA$. Since $\frac{N}{d_{i,v}}D_{i,v}\sim NA$, there exists some hyperplane $H_{i,v}\in\mathbb{P}^{M}$ defined over $k$ such that $\frac{N}{d_{i,v}}D_{i,v}=\phi^{*}H_{i,v}$. For $P\in X(k)\setminus\operatorname{Supp}D_{i,v}$, by functoriality and additivity of Weil functions, we have (up to $O(1)$)
$\log\boxdot_{H_{i,v},v}(\phi(P))=\log\boxdot_{\frac{N}{d_{i,v}}D_{i,v},v}(P)=N\log\boxdot^{\frac{1}{d_{i,v}}}_{D_{i,v},v}(P).$
We also find $h(\phi(P))=Nh_{A}(P)+O(1)$ for $P\in X(k)$. Plugging these identities into (13) in the case of hyperplanes with $c_{i,v}=1$, we can get the desired result.
Ultimately, we demonstrate that the statement remains valid even if we replace linear equivalence with numerical equivalence. This conclusion is supported by the following result due to Matsusaka [Mat58] in algebraic geometry:
Lemma 3.3 (Matsusaka).
Let $A$ be an ample Cartier divisor on a projective variety $X$. For any Cartier divisor $D$, there exists a positive integer $N_{0}$ satisfying for all $N\geq N_{0}$, then $D\equiv NA$, $D$ is very ample.
Now we can form the following family of divisors $Q_{i,v}$ to replace $D_{i,v}$.
Let $d$ be the least common multiple of the integers $d_{i,v}$. By replacing $D_{i,v}$ with $\frac{dD_{i,v}}{d_{i,v}}$ for all $i$, and $A$ with $dA$, and leveraging the additivity of Weil functions and heights (up to bounded functions), we observe that it is sufficient to prove the case where $d_{i,v}=1$ for all $i$ and all $v\in S$, i.e., $D_{i,v}\equiv A$ for all $i$ and all $v\in S$. Let $N_{0}$ be the integer in lemma 3.3 for $A$. By standard properties of Weil functions and heights, there exists an integer $N>N_{0}$ such that
(14) $\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot^{\frac{N_{0}}{N}}_{D_{i,v},v}(P)<\epsilon h_{A}(P)+O(1)$
for all $P\in X(k)\setminus\cup_{i,v}\operatorname{Supp}D_{i,v}$. By the selection of $N_{0}$, it follows that $NA-(N-N_{0})D_{i,v}$ is very ample for all $i$ and $v\in S$. According to this fact, there exist effective divisors $E_{i,v}$ such that $NA\sim(N-N_{0})D_{i,v}+E_{i,v}$, and for all $v\in S$, the divisors $(N-N_{0})D_{1,v}+E_{1,v},\ldots,(N-N_{0})D_{q,v}+E_{q,v}$ and $D_{1,v},\ldots,D_{q,v}$ have the same characteristic of position. We now apply (13) to the linearly equivalent divisors $Q_{i,v}=(N-N_{0})D_{i,v}+E_{i,v}$, $i=1,\ldots,q$, $v\in S$, and $NA$. We obtain
$\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot_{(N-N_{0})D_{i,v}+E_{i,v},v}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{NA}(P)$
for all $P\in X(k)\setminus Z$ for some proper Zariski-closed subset $Z$ of $X$ containing the supports of all $E_{i,v}$, $v\in S$, $i=1,\ldots,q$. Since $\log\boxdot_{E_{i,v},v}$ is bounded from below outside of the support of $E_{i,v}$, we obtain
$\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot^{\left(1-\frac{N_{0}}{N}\right)}_{D_{i,v},v}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{A}(P)+O(1)$
for all $P\in X(k)\setminus Z$. Using (14) we find
(15) $\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot_{D_{i,v},v}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+2\epsilon\right)h_{A}(P)+O(1)$
for all $P\in X(k)\setminus Z$. Given that there are only finitely many $k$-rational points with bounded height (with respect to $A$), by adding a finite number of points to $Z$, we can ensure that equation (15) holds with $2\epsilon$ replaced by $3\epsilon$ without the $O(1)$ term. So far, we have fully completed the proof of Part I.
Part II: An inequality within parameters and controlling the dimension of the exceptional set $Z$.
In consideration of the numerically equivalent ample divisors, now we begin prove(i.e. proof of the relatively complex side of the index in the inequality of [Lev14])
$C(\delta,l,m,n)\leq B(l,\delta m,\delta n)\leq\max_{l\leq j\leq\delta n}A(\delta m,j).$
Consider the quantities and objects $c,l,m,n,k,S,X,A,D_{1,v},\ldots,D_{q,v}$(where $q=m+1$), and $D$ as previously defined. Let $Z$ denote the exceptional set referred to in equation (6) with $\delta$ be a positive integer. Assume that the dimension of $Z$ is $j\geq l$. If $\delta=1$, we let $W$ be an irreducible component of $Z$ (considered over the algebraic closure $\overline{k}$ ) such that $\dim W=j$. By extending the field $k$ if necessary, we may assume that $W$ is defined over $k$. The restricted divisors $D_{1,v}|_{W},\ldots,D_{q,v}|_{W}$ are in $m$-subgeneral position on $W$, and the restricted ample divisor $A|_{W}$ remains ample on $W$. By properties of Weil functions and the setup of $Z$ and $W$, for any $\epsilon>0$ the set of $P\in W(k)\setminus\operatorname{Supp}D$ such that
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot^{\epsilon_{D_{{i,v}|W}}(A)}_{D_{{i,v}|W},w}(P)>(c-\epsilon)h_{A|_{W}}(P)$
is Zariski-dense in $W$. It shows that $A(m,j)\geq c$ which implies that $B(l,m,n)\leq\max_{l\leq j\leq n}A(m,j)$.
Let $\epsilon>0$. Form the set $Z$ defined to be the Zariski-closure of the set
$\left\{P\in X(\overline{k})\setminus\operatorname{Supp}D\mid\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot^{\epsilon_{D_{i,v}}(A)}_{D_{i,v},w}(P)>(B(l,\delta m,\delta n)+\epsilon)h_{A}(P)\right\}$
for $[k(P):k]=\delta$ (we assume without of loss of generality that $B(l,\delta m,\delta n)$ is finite in our following proof). It suffices to show that $\dim Z<l$, which is equivalent to prove $C(\delta,l,m,n)\leq B(l,\delta m,\delta n)$.
To simplify the proof and without loss of generality, replacing $\left(\prod_{j\neq i}\epsilon_{D_{i,v}}(A)\right)D_{i,v}$ by $D_{i,v}$ and $\left(\prod_{j=1}^{q}\epsilon_{D_{i,v}}(A)\right)A$ by $A$, it suffices to prove the case where $\epsilon_{D_{i,v}}(A)=1$ for all $i$.
[…]
(i) Notation and Definitions: Let $X^{\delta}$ denote the $\delta$th power of $X$, which is the Cartesian product of $X$ with itself $\delta$ times. Let $X^{(\delta)}$ denote the $\delta$-th symmetric power of $X$, defined as the quotient $X^{\delta}/S_{\delta}$, where $S_{\delta}$ is the symmetric group on $\delta$ elements acting by permuting the coordinates of $X^{\delta}$. Let $\phi:X^{\delta}\to X^{(\delta)}$ be the natural projection map that takes a point in $X^{\delta}$ to its equivalence class in $X^{(\delta)}$.
(ii) Points and Conjugates: For a point $P\in X(\overline{k})$ (where $\overline{k}$ is the algebraic closure of $k$) satisfying $[k(P):k]=\delta$, let $P=P_{1},P_{2},\ldots,P_{\delta}$ denote the $\delta$ distinct conjugates of $P$ over $k$. Define the map $\psi:X(\overline{k})\to X^{\delta}$ by $\psi(P)=(P_{1},P_{2},\ldots,P_{\delta})$.
(iii) Projection Maps: Let $\pi_{i}:X^{\delta}\to X$ denote the $i$th projection map, which projects a point in $X^{\delta}$ onto its $i$-th coordinate.
(iv) Push-forward and Pullback: Define $D^{*}_{i,v}=\phi_{*}\pi_{1}^{*}(D_{i,v})$, where $D_{i,v}$ are divisors on $X$. Define $D^{*}=\sum_{v\in S}\sum_{i=1}^{q}D^{*}_{i,v}$. Define $A^{*}=\phi_{*}\pi_{1}^{*}(A)$, where $A$ is an ample divisor on $X$.
(v) Properties of Divisors: Note that $\phi^{*}\phi_{*}\pi_{1}^{*}D_{i,v}=\sum_{j=1}^{\delta}\pi_{j}^{*}D_{i,v}$. Since the $D_{i,v}$ are in $m$-subgeneral position, the $D^{*}_{i,v}$ are in $m\delta$-subgeneral position. If $D_{i,v}\equiv D_{j,v}\equiv A$ for all $i$ and $j$ , then $D^{*}_{i,v}\equiv D^{*}_{j,v}\equiv A^{*}$ for all $i$ and $j$.
(vi) Ampleness and Effectiveness: Since $A$ is ample, $A^{*}$ is also ample. Consequently, $D^{*}_{i,v}$ is effective and ample for all $i$ and $v$.
(vii) Rational Points: If $P\in X(\overline{k})$ and $[k(P):k]=\delta$, then the point $\phi(\psi(P))\in X^{(\delta)}$ is $k$-rational.
Thus by the definition of $B(l,\delta m,\delta n)$, we have the inequality
(16) $\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D^{*}_{i,v},w}(P)<(B(l,\delta m,\delta n)+\epsilon)h_{A^{*}}(P)$
$\quad\forall P\in X^{(\delta)}(k)\setminus(Z^{\prime}\cup\operatorname{Supp}D^{*}),$
for some Zariski-closed subset $Z^{\prime}$ of $X^{(\delta)}$ satisfying $\dim Z^{\prime}\leq l-1$. For $P\in X(\overline{k})\setminus\operatorname{Supp}D$, $[k(P):k]=\delta$, we have, up to $O(1)$,
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D^{*}_{i,v},w}(\phi(\psi(P)))$
[…]
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{\phi_{*}\pi_{1}^{*}D_{i,v},w}(\phi(\psi(P)))$
[…]
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{\phi^{*}\phi_{*}\pi_{1}^{*}D_{i,v},w}(\psi(P))=\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{\sum_{j=1}^{\delta}\pi_{j}^{*}D_{i,v},w}(\psi(P))$
[…]
$\sum_{j=1}^{\delta}\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{\pi_{j}^{*}D_{i,v},w}(\psi(P))$
[…]
$\sum_{j=1}^{\delta}\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D_{i,v},w}(\pi_{j}\psi(P))=\sum_{j=1}^{\delta}\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D_{i,v},w}(P_{j})$
[…]
$\delta\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D_{i,v},w}(P).$
We also have $h_{A}^{*}(\phi(\psi(P)))=\delta h_{A}(P)+O(1)$. Consider $Z^{*}\subset Z(\overline{k})$ is a Zariski-dense set of points in $Z$ satisfying $[k(P):k]=\delta$. It follows that for an appropriate choice of the functions,
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D^{*}_{i,v},w}(\phi(\psi(P)))>(B(l,\delta m,\delta n)+\epsilon)h_{A^{*}}(\phi(\psi(P)))$
for all $P\in Z^{*}$. From (16), we have $\dim\phi(\psi(Z^{*}))\leq l-1$. Since $\phi$ is a finite map and $\pi_{1}(\psi(Z^{*}))=Z^{*}$, $\dim\phi(\psi(Z^{*}))\geq\dim Z^{*}=\dim Z$. Thus $\dim Z\leq l-1$, as desired. It also shows that $C(\delta,l,m,n)\leq B(l,\delta m,\delta n)$. Based on the above construction and discussion, we can also obtain $B(l,\delta m,\delta n)\leq\max_{l\leq j\leq\delta n}A(\delta m,j)$.
Part III: Part II for algebraic points of bounded degree.
Set $c_{i,v}=1$ for all $i$ and $v$(that’s the point we focus on)in Part I’s result and integrate with the part II, the factor on the right side as stated above modify into $a_{\Delta,\delta}$, where $a_{\Delta,\delta}$ here is defined by
$\left[(n+1)\max_{v\in S}\delta_{\mathcal{D}_{v}}(i.e-\Delta)\right]_{w.r.t-\delta}$
while
$\delta_{\mathcal{D}_{v}}=\max_{I\subset\{1,\ldots,q\}}\max\left\{1,\frac{\#I}{\operatorname{codim}\cap_{i\in I}\operatorname{Supp}D_{i}}\right\}=\max\left\{1,\max_{\varnothing\subsetneq W\subsetneq X}\frac{\alpha_{v}(W)}{\operatorname{codim}W}\right\},$
where we set $\operatorname{codim}\varnothing=\infty$.
To explain clearly, we are here to show that the above two definitions are equivalent(although we have already mention in introduction).
Assume that $\tilde{W}$ is a subvariety of $X$ where the ratio $\frac{\alpha_{v}(W)}{\operatorname{codim}W}$ achieves its maximum value precisely when $W=\tilde{W}$. By reordering if necessary, we can assume that $\tilde{W}\subset D_{i}$ for $i=1,\dots,\alpha_{v}(\tilde{W})$. Define $W^{\prime}=\bigcap_{i=1}^{\alpha_{v}(\tilde{W})}D_{i}$. It is evident that $\tilde{W}\subset W^{\prime}$, which implies that $\operatorname{codim}\tilde{W}\geq\operatorname{codim}W^{\prime}$. Additionally, since $\tilde{W}\nsubseteq D_{i}$ for all $i>\alpha_{v}(\tilde{W})$, it follows that $W^{\prime}\nsubseteq D_{i}$ for all $i>\alpha_{v}(\tilde{W})$. Therefore, we conclude that $\alpha_{v}(\tilde{W})=\alpha_{v}(W^{\prime})$. Thus, we have
$\frac{\alpha_{v}(W^{\prime})}{\operatorname{codim}W^{\prime}}\geq\frac{\alpha_{v}(\tilde{W})}{\operatorname{codim}\tilde{W}}.$
By our assumption for $\tilde{W}$, we get
$\frac{\alpha_{v}(W^{\prime})}{\operatorname{codim}W^{\prime}}=\frac{\alpha_{v}(\tilde{W})}{\operatorname{codim}\tilde{W}}.$
This means we only need to consider those $W$ which are the intersections of some $D_{i}$’s, and our claim follows from this observation.
Combined with the inequality in Part II we have,
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $D_{1,v},\ldots,D_{q,v}(q=m+1)$ be effective ample Cartier divisors of $X$, defined over $k$, set Seshadri constants $\epsilon_{D_{i,v}}(A)$ construct positive integers $d_{i,v}$ such that $D_{i,v}\equiv d_{i,v}A$ for all $i$ and for all $v\in S$, then
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot^{\epsilon_{D_{i,v}}(A)}_{D_{i,v},w}(P)<\left(a_{\Delta,\delta}+\epsilon\right)h_{A}(P)+O(1)$
for all points $P\in X(\bar{k})\setminus\cup_{i,v}\operatorname{Supp}D_{i,v}$, as desired.
At this point, we have proven the theorem 1.7 with the weights index $c_{i,v}=1$. In comparison with theorem 1.4, due to other cases special nature, there is no unified trivial representation, so we use the form of our main result for resolution.
Using the definition if the $D_{i,v}$ are taken to be effective(Cartier) divisors in $m$-subgeneral position, then
$\Delta=\max_{v\in S}\delta_{\mathcal{D}_{v}}\leq m-n+1$
The inequality in Part II shows that,
$C(\delta,1,m,n)\leq B(1,\delta m,\delta n)\leq\max_{1\leq j\leq\delta n}A(\delta m,j)\leq\max_{1\leq j\leq\delta n}\left(\delta m-j+1\right)(j+1)=a_{\Delta,\delta}$
Consider $\left(\delta m-j+1\right)(j+1)$ as a quasi-quadratic function, based on the most basic knowledge and the value of variable must be integer, we have
(i)For $m>2n$,
$a_{\Delta,\delta}=n({\delta^{2}}m-{\delta^{2}}n)+{\delta}m+1$
(ii)For $m\leq 2n$, in this case, $j=\left[\frac{\delta m}{2}\right]_{Int}$(notation in introduction), so we have
$a_{\Delta,\delta}=-j^{2}+\delta mj+{\delta}m+1$
and if $D_{i,v}$ are divisors in $m$-subgeneral position with index $\kappa$, then
$\Delta\leq\frac{m-n}{\kappa}+1$
thus,
$C(\delta,1,m,n)\leq B(1,\delta m,\delta n)\leq\max_{1\leq j\leq\delta n}A(\delta m,j)\leq\max_{1\leq j\leq\delta n}\left(\frac{\delta m-j}{\kappa}+1\right)(j+1).$
The same operation as above,
(i)For $m>\lceil 2n+\frac{1-\kappa}{\delta}\rceil$,
$a_{\Delta,\delta}=\frac{n}{\kappa}\left({\delta^{2}}m-{\delta^{2}}n+\delta(\kappa-1)\right)+\frac{{\delta}m}{\kappa}+1$
(ii)For $m\leq\lfloor 2n+\frac{1-\kappa}{\delta}\rfloor$, as above, $j=\left[\frac{\delta m+(\kappa-1)}{2}\right]_{Int}$ in this case, then
$a_{\Delta,\delta}=-\frac{j^{2}}{\kappa}+(\frac{\delta m}{\kappa}-\frac{1}{\kappa}+1)j+\frac{{\delta}m}{\kappa}+1$
thus we completely end our proof. ∎