### v5

# A generalized Schmidt’s subspace theorem for algebraic points of bounded degree

Abstract.
Schmidt’s subspace theorem in terms of Seshadri constants for closed subschemes in subgeneral position has been already developed sharply. We derive our theorem for numerically equivalent ample divisors by dint of the above theory step by step, providing a joint weights version of generalized Schmidt’s theorem for algebraic points of bounded degree.
Key words and phrases:
Schmidt’s subspace theorem, Roth’s theorem, Diophantine approximation, Divisors, Seshadri constants, Algebraic point
Mathematics Subject Classification (2020): 11J87, 14G40, 11J25, 11J68, 11D75, 37P30.
1. Introduction
1.1. Glorious work for Schmidt’s subspace theorem
Schmidt’s subspace theorem is a fundamental component in the realm of higher-dimensional Diophantine approximation. Over the years, it has been extended and generalized by numerous researchers. Notable contributions have been made by Evertse and Ferretti (as seen in [EF08]), Corvaja and Zannier (as seen in [CZ04a]), and Ru and Vojta [RV20]. A recent version of this theorem, which involves closed subschemes in general position and their Seshadri constants relative to a given ample divisor $A$, has been established. This version builds upon the work of Ru and Wang [RW22] and Vojta [Voj20], who have explored the use of beta constants as an alternative to Seshadri constants. This particular formulation was proven in the recent paper [HL21].
Theorem 1.1 ([HL21, Theorem 1.3]).
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$. Let $S$ be a finite set of places of $k$. For each $v\in S$, let $Y_{0,v},\ldots,Y_{n,v}$ be closed subschemes of $X$ in general position, defined over $k$. Let $A$ be an ample Cartier divisor on $X$. Then there exists a proper Zariski-closed subset $Z\subset X$ such that for all points $P\in X(k)\setminus Z$ and $\epsilon>0$,
$\sum_{v\in S}\sum_{i=0}^{n}\epsilon_{Y_{i,v}}(A)\lambda_{Y_{i,v},v}(P)<(n+1+\epsilon)h_{A}(P).$
More recently, in [Qua22a], Quang introduced the concept of a distributive constant $\Delta$ for a family of divisors, and later in the [Qua22b], Quang also expanded this notable notion to encompass closed subschemes:
Definition 1.2.
For a family of closed subschemes (similar for divisors) $\mathcal{Y}=\{Y_{1},\ldots,Y_{q}\}$ defined on a projective variety $X$, the distributive constant is given by
$\delta_{\mathcal{Y}}=\max_{J\subset\{1,\ldots,q\}}\max\left\{1,\frac{\#J}{\operatorname{codim}\cap_{j\in J}\operatorname{Supp}Y_{j}}\right\},$
where we set $\operatorname{codim}\varnothing=\infty$.
This raises an important generalized theorem:
Theorem 1.3 (Quang [Qua22b]).
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $Y_{1,v},\ldots,Y_{q,v}$ be closed subschemes of $X$, defined over $k$. Let $\Delta=\max_{v\in S}\delta_{\mathcal{Y}_{v}}$ be the maximum of the distributive constants of $\mathcal{Y}_{v}=\{Y_{1,v},\ldots,Y_{q,v}\}$. Let $A$ be an ample Cartier divisor on $X$. Then there exists a proper Zariski closed set $Z$ of $X$ such that
$\sum_{v\in S}\sum_{i=1}^{q}\epsilon_{Y_{i,v}}(A)\lambda_{Y_{i,v},v}(P)<\left(\Delta(n+1)+\epsilon\right)h_{A}(P)$
for all points $P\in X(k)\setminus Z$ and $\epsilon>0$.
As discussed in [Qua22a],
$\Delta\leq m-n+1$
if the $Y_{i,v}$ are taken to be closed subschemes in $m$-subgeneral position.
$\Delta\leq\frac{m-n}{\kappa}+1$
if the $Y_{i,v}$ are taken to be closed subschemes in $m$-subgeneral position with index $\kappa$.
Therefore, Theorem 1.3 generalizes and encompasses the prior results concerning closed subschemes in $m$-subgeneral position and its variations.
In addition, it is a breakthrough that in the latest paper [HL25], the authors have extend Theorem 1.3 to the following Schmidt-Nochka-type Theorem for weighted sums involving arbitrary closed subschemes.
Theorem 1.4 ([HL25], Theorem 1.2).
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $Y_{1,v},\ldots,Y_{q,v}$ be closed subschemes of $X$ defined over $k$, and let $c_{1,v},\ldots,c_{q,v}$ be nonnegative real numbers. For a closed subset $W\subset X$ and $v\in S$, let
$\alpha_{v}(W)=\sum_{\begin{subarray}{c}i\\ W\subset\operatorname{Supp}Y_{i,v}\end{subarray}}c_{i,v}.$
Let $A$ be an ample Cartier divisor on $X$. Then there exists a proper Zariski closed subset $Z$ of $X$ and $\epsilon>0$ such that
$\sum_{v\in S}\sum_{i=1}^{q}c_{i,v}\epsilon_{Y_{i,v}}(A)\lambda_{Y_{i,v},v}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{A}(P)$
for all points $P\in X(k)\setminus Z$.
The application of weights in the inequality described above is entirely innovative and substantial, which provides convenience for addressing a variety of situations.
In comparison with Theorem 1.3, under the assumption $c_{i,v}=1$ for all $i$ and $v$, then we assert that
$\delta_{\mathcal{Y}_{v}}=\max\left\{1,\max_{\varnothing\subsetneq W\subsetneq X}\frac{\alpha_{v}(W)}{\operatorname{codim}W}\right\}.$
This conclusion stems directly from the definitions and a straightforward observation (which we will elaborate on in our proof section-Part III, in light of the detailed explanation in [CY25]). When taking the maximum over $W$, it suffices to consider the cases where $W$ is the intersection of a subset of closed subschemes of $\mathcal{Y}_{v}$.
Overall, as stated in [HL25], Theorem 1.4 can be regarded as a weighted extension of Theorem 1.3 for algebraic points. Specifically, in certain cases, Theorem 1.4 improves upon the coefficient of the height compared to Quang’s theorem, primarily due to the presence of the maximum with 1 in Quang’s definition of $\delta_{\mathcal{Y}}$. And it significantly prune the discussion requirements for the condition of position. Moreover, this theorem represents the most powerful generalization available to date.
1.2. Constructive results in Schmidt’s theorem for algebraic points of bounded degree
Previous work studying Schmidt’s theorem (for numerically equivalent ample divisors) for algebraic points of bounded degree proved the following theorem (reformulate in the form of hypersurfaces in $\mathbb{P}^{N}$),
Theorem 1.5 ([Lev14], Theorem 1.7).
Let $X$ be a projective subvariety of $\mathbb{P}^{N}$ with dimension $n\geq 1$ defined over a number field $k$. Let $S$ be a finite set of places of $k$. For $v\in S$, let $H_{0,v},\ldots,H_{n,v}\subset\mathbb{P}^{N}$ be hypersurfaces over $k$ in n-subgeneral position (general position). Let $\delta\geq 1$ be an integer and $\epsilon>0$. Then the inequality
(1) $\sum_{v\in S}\sum_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\sum_{i=0}^{n}\frac{\lambda_{H_{i,v},w}(P)}{\deg H_{i,v}}<\left((\delta n)^{2}\left(\frac{\delta n-1}{2\delta n-3}\right)+\epsilon\right)h(P)$
holds for all but finitely many points $P\in X(\overline{k})\setminus\cup_{v\in S}\cup_{i=0}^{n}H_{i,v}$ satisfying $[k(P):k]\leq\delta$.
Set $X=\mathbb{P}^{n}=\mathbb{P}^{N}$, while $H_{i,v}\subset X$ are hyperplanes, Schlickewei ([Sch03]) proposed that there exists a constant $c(\delta,n)$, which depends solely on $\delta$ and $n$, and a finite union of hyperplanes $Z\subset\mathbb{P}^{n}$ such that the left-hand side of inequality (1) is bounded by $(c(\delta,n)+\epsilon)h(P)$ for all points $P\in\mathbb{P}^{n}(\overline{k})\setminus Z$ satisfying $[k(P):k]\leq\delta$. Consequently, Theorem 1.5 confirms Schlickewei’s conjecture as a special case.
Combined with the inequality that we will prove in great detail, if $\delta m\geq 2$, we have a more general theorem (Theorem 1.6) proved in [Lev14], i.e. own a factor of
$C(\delta,1,m,n)\leq\max_{1\leq j\leq\delta n}A(\delta m,j)\leq\max_{1\leq j\leq\delta n}\frac{\delta m(\delta m-1)(j+1)}{\delta m+j-2}$
$\leq$ $\frac{\delta m(\delta m-1)(\delta n+1)}{\delta m+\delta n-2}.$
Explicitly, this provides an adaptation of Schmidt’s theorem for algebraic points.
Theorem 1.6 ([Lev14], Theorem 5.2).
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$. Let $S$ be a finite set of places of $k$. Let $D_{0,v},\ldots,D_{m,v}$ be effective ample Cartier divisors on $X$ in $m$-subgeneral position for each $v\in S$. Suppose that there exists a Cartier divisor $A$ on $X$ defined over $k$, and positive integers $d_{i,v}$ such that $D_{i,v}\equiv d_{i,v}A$ for all $i$ and for all $v\in S$. Let $\delta$ be a positive integer with $\delta m\geq 2$. Then the inequality
$\sum_{v\in S}\sum_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\sum_{i=0}^{m}\frac{\lambda_{D_{i,v},w}(P)}{d_{i,v}}<\left(\frac{\delta m(\delta m-1)(\delta n+1)}{\delta m+\delta n-2}+\epsilon\right)h_{A}(P)+O(1)$
holds for all points $P\in X(\overline{k})\setminus\cup_{i,v}\operatorname{Supp}D_{i,v}$ satisfying $[k(P):k]\leq\delta$ and $\epsilon>0$.
If $m=n\geq 2$, then using $A(\delta n,\delta n)\leq\delta n+1$ gives the following slightly better bound (which is essentially quadratic in $\delta n$) compared with the previous relevant conclusions in higher dimension and for larger $\delta$:
$C(\delta,1,n,n)\leq\max_{1\leq j\leq\delta n}A(\delta n,j)\leq(\delta n)^{2}\frac{\delta n-1}{2\delta n-3}.$
Comprehensively, inspired by the newest work in Schmidt’s subspace theorem, the following theorem is our main result:
Theorem 1.7.
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $D_{1,v},\ldots,D_{q,v}$ ($q=m+1$) be effective (ample) Cartier divisors of $X$ defined over $k$ in $m$-subgeneral position, set Seshadri constants $\epsilon_{D_{i,v}}(A)$ corresponding to positive integers $d_{i,v}$ such that $D_{i,v}\equiv d_{i,v}A$ for all $i$ and for all $v\in S$, where $A$ is an ample Cartier divisor defined on $X$. For a closed subset $W\subset X$ and $v\in S$, let
$\alpha_{v}(W)=\#\left\{i\mid W\subset\operatorname{Supp}D_{i,v}\right\}.$
Then there exists $\epsilon>0$ and $\delta$ be a positive integer, such that
$\sum_{v\in S}\sum_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\sum_{i=1}^{q}\epsilon_{D_{i,v}}(A)\lambda_{D_{i,v},w}(P)<\left(a_{\delta}+\epsilon\right)h_{A}(P)+O(1)$
where
$a_{\delta}=\left[(n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}\right]_{w.r.t-\delta}$
holds for all points $P\in X(\overline{k})\setminus\cup_{i,v}\operatorname{Supp}D_{i,v}$ satisfying $[k(P):k]\leq\delta$.
$[*]_{w.r.t-\delta}$ here is a consequence of an operation which invoking the method that in our proof section and more complementary details can be seen in Appendix part. In view of the particularity of various situations, we will systematically illustrate and list the results in the following remark and uniformly use $m$-subgeneral position and $c_{i,v}=1$ as the main research object to get some slightly more general solution (e.g., with index $\kappa$), which is equivalent to a sharp integration and extension of Levin’s results in [Lev14]. Also we will explain the meaning of $\epsilon_{D_{i,v}}(A)$ corresponding to positive integers $d_{i,v}$ in our proof section-Part I, presented in $\mathbf{Perspective}$-ii. We can consider this as a generalized Seshadri constant setup, which is broader than the original definition but has imposed auxiliary assignment constraints.
Remark 1.8.
Let $X\subset\mathbb{P}^{N}(k)$ be a projective variety of dimension $n$ defined over a number field $k$, if $D$ is a hypersurface in $\mathbb{P}^{N}(k)$ defined by a homogeneous polynomial $f\in k[x_{0},\ldots,x_{N}]$ of degree $d$, for $v\in M_{k}$, we let $|f|_{v}$ denote the maximum of the absolute values of the coefficients of $f$ with respect to $v$. We define $\|f\|_{v}$ similarly, then a Weil function for $D$ is given by
$\lambda_{D,v}(P)=\log\frac{\|f\|_{v}\max_{i}\|x_{i}\|_{v}^{d}}{\|f(P)\|_{v}}.$
so for $\mathcal{D}_{v}=\{D_{1,v},\ldots,D_{q,v}\}(q=m+1)$ be hypersurfaces, we have,
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{I=1}^{q}\left(\frac{\|f_{I}\|_{v,w}\max_{0\leq i\leq N}\|x_{I_{i}}\|^{d_{I}}_{v,w}}{\|f_{I}(P)\|_{v,w}}\right)^{\epsilon_{D_{I,v}}(A)}<\left(a_{\delta}+\epsilon\right)h_{A}(P)+O(1)$
holds for all points $P\in X(\bar{k})\setminus\cup_{I,v}\operatorname{Supp}D_{I,v}$ satisfying $[k(P):k]\leq\delta$. (In fact, with the assumption as above, $\epsilon_{D_{I,v}}(A)=\frac{1}{d_{I,v}}$ in this situation.)
By our assumption and discussion in proof section, we slightly weaken the intensity of the factor, thereby obtaining a result that is intuitive and general, we modify $a_{\delta}$ into $a_{\Delta,\delta}$, where $a_{\Delta,\delta}$ here is defined by:
$\left[(n+1)\max_{v\in S}\delta_{\mathcal{D}_{v}}(i.e.-\Delta)\right]_{w.r.t-\delta}$
Then we get the following main results for $m$-subgeneral position:
(i)For $m>2n$,
$a_{\Delta,\delta}=n({\delta^{2}}m-{\delta^{2}}n)+{\delta}m+1.$
(ii)For $n\leq m\leq 2n$, set $[*]_{Int}$ as a rounding symbol, taking the nearest positive integer to * (or just itself if * is already an integer. When * is exactly the symmetry axis, then two symmetrical integers can meet the criteria), thus for $j$ ($1\leq j\leq\delta n$) defined as above, we have the maximum when $j=\left[\frac{\delta m}{2}\right]_{Int}\Leftrightarrow i$, which shows that:
$a_{\Delta,\delta}=-i^{2}+\delta m\cdot i+{\delta}m+1.$
This improved an earlier result in [Lev14] which had a factor of
$\frac{\delta m(\delta m-1)(\delta n+1)}{\delta m+\delta n-2}+\epsilon$
on the right-hand side. Furthermore, for more general case $m$-subgeneral position with index $\kappa$, we can definitely guarantee that:
(i)For $m\geq\lceil 2n+\frac{1-\kappa}{\delta}\rceil$,
$a_{\Delta,\delta}=\frac{n}{\kappa}\left({\delta^{2}}m-{\delta^{2}}n+\delta(\kappa-1)\right)+\frac{{\delta}m}{\kappa}+1.$
(ii)For $n\leq m\leq\lfloor 2n+\frac{1-\kappa}{\delta}\rfloor$, in the same way, $j=\left[\frac{\delta m+(\kappa-1)}{2}\right]_{Int}\Leftrightarrow i$,
$a_{\Delta,\delta}=-\frac{i^{2}}{\kappa}+\left(\frac{\delta m+(\kappa-1)}{\kappa}\right)i+\frac{{\delta}m}{\kappa}+1.$
In this generalized case, due to the limitation of expression of $a_{\delta}$, the factor $a_{\Delta,\delta}$ is more representative in contrast to $a_{\delta}$ and also yields an insurmountable result.
The structure of this paper is as follows. In Section 2, we recall some basic definitions and introduce some notions. In Section 3, we will divide the problem into three parts, initially provide a rigorous proof of Schmidt’s subspace theorem for effective Cartier divisors based on the idea in [HL25] and extent it to numerically equivalent ample divisors in two different ways, and then handle the issue of algebraic points of bounded degree.
2. Definitions and preliminaries
Throughout this paper, we denote $k$ as a number field while $\overline{k}$ is algebraic closure of $k$, and $\mathcal{O}_{k}$ as the ring of integers of $k$. We define $M_{k}$ as a canonical set of places (or absolute values) of $k$, which includes: (i) One place for each prime ideal $\mathfrak{p}$ of $\mathcal{O}_{k}$, (ii) One place for each real embedding $\sigma:k\to\mathbb{R}$, (iii) One place for each pair of conjugate embeddings $\sigma,\overline{\sigma}:k\to\mathbb{C}$.
Given a finite set $S$ of places of $k$ that contains all the archimedean places, we denote $\mathcal{O}_{k,S}$ as the ring of $S$-integers of $k$, $\mathcal{O}_{k,S}^{*}$ as the group of $S$-units of $k$. For a place $v$ of $k$ and a place $w$ of a field extension $L$ of $k$, we say that $w$ lies above $v$ (denoted $w|v$) if $w$ and $v$ induce the same topology on $k$. We normalize the absolute values as follows: $|p|_{v}=\frac{1}{p}$ if $v$ corresponds to a prime ideal $\mathfrak{p}$ lying above a rational prime $p$, $|x|_{v}=|\sigma(x)|$ if $v$ corresponds to an embedding $\sigma$. For each $v\in M_{k}$, we let $k_{v}$ denote the completion of $k$ with respect to $v$. We set
$\|x\|_{v}=|x|_{v}^{[k_{v}:\mathbb{Q}_{v}]/[k:\mathbb{Q}]}.$
A fundamental equation is the product formula
$\prod_{v\in M_{k}}\|x\|_{v}=1,$
which holds for all $x\in k^{*}$.
We start by providing a concise overview of the essential properties of local and global height functions, primarily following the reference [Sil87] and [HL25]. Subsequently, we will introduce the concept of subgeneral position and Seshadri constants, which are utilized in this paper.
For more refined versions and more details about local (global) height functions can be found in [HL25], we will only provide a brief explanation here.
2.1. Local height functions
Let $Y$ be a closed subscheme of a projective variety $X$, both defined over a number field $k$. For any place $v$ of $k$, one can associate a local height function (or Weil function) $\lambda_{Y,v}:X(k)\setminus Y\to\mathbb{R}$, up to a bounded term $O(1)$. This function measures the $v$-adic distance of a point in $X(k)\setminus Y$ to the closed subscheme $Y$, taking large values when the point is close to $Y$. In particular, $\lambda_{Y,v}$ is bounded from below for all $P\in X(k)\setminus Y$. It satisfy the following properties: if $Y$ and $Z$ are two closed subschemes of $X$, defined over $k$, and $v$ is a place of $k$, then up to $O(1)$,
$\lambda_{Y\cap Z,v}=\min\{\lambda_{Y,v},\lambda_{Z,v}\},\lambda_{Y+Z,v}=\lambda_{Y,v}+\lambda_{Z,v},\lambda_{Y,v}\leq\lambda_{Z,v},\text{ if }Y\subset Z.$
It is worth highlighting that by setting $Y=Z$, we obtain the following equality
$\lambda_{2Y,v}=\lambda_{Y+Y,v}=\lambda_{Y,v}+\lambda_{Y,v}=2\lambda_{Y,v},$
where $2Y$ denotes the closed subscheme corresponding to the ideal sheaf $\mathcal{I}_{Y}^{2}$. Furthermore, it is generally true that if $Y_{1},\ldots,Y_{r}$ are closed subschemes of $X$ and $c_{1},\ldots,c_{r}$ are positive integers, we have
(2) $\lambda_{c_{1}Y_{1}\cap\ldots\cap c_{r}Y_{r},v}$ $=\min\{c_{1}\lambda_{Y_{1},v},\ldots,c_{r}\lambda_{Y_{r},v}\},$
(3) $\lambda_{c_{1}Y_{1}+\cdots+c_{r}Y_{r},v}$ $=c_{1}\lambda_{Y_{1},v}+\cdots+c_{r}\lambda_{Y_{r},v}.$
If $\phi:W\to X$ is a morphism of projective varieties with $\phi(W)\not\subset Y$, then up to $O(1)$,
$\lambda_{Y,v}(\phi(P))=\lambda_{\phi^{*}Y,v}(P),$
$\quad\forall P\in W(k)\setminus\phi^{*}Y.$
The operations $Y\cap Z$, $Y+Z$, $Y\subset Z$, and $\phi^{*}Y$ are defined in terms of the associated ideal sheaves. Specifically, it is a subtle point that if $Y$ is associated with the ideal sheaf $\mathcal{I}_{Y}$, then $\phi^{*}Y$ represents the closed subscheme corresponding to the inverse image ideal sheaf $\phi^{-1}\mathcal{I}_{Y}\cdot\mathcal{O}_{W}$.
If $Y=D$ is an effective Cartier divisor (which we often identify with the associated closed subscheme according to the Lemma 2.2 in [Sil87]), these height functions coincide with the usual height functions associated to divisors.
2.2. Global height functions
As our research focus, global height functions of Cartier divisors, especially ample Cartier divisors, play a crucial role. Same definitions apply for $Y$ and $X$, a global height function $h_{Y}:X(k)\setminus Y\to\mathbb{R}$, can be associated as follows,
$h_{Y}(P)=\sum_{v\in M_{k}}\lambda_{Y,v}(P).$
For $v\in M_{k}$, the definition of $\lambda_{Y,v}$ can be chosen such that the collection $(\lambda_{Y,v})_{v\in M_{k}}$ is well-defined up to an $M_{k}$-constant (a sequence of real numbers $(c_{v})_{v\in M_{K}}$ where $c_{v}=0$ for all but finitely many $v$). As a result, $h_{Y}$ is well-defined up to $O(1)$.
These functions possess properties similar to those described earlier for local height functions, with one notable exception: the property above is modified to $h_{Y\cap Z}\leq\min\{h_{Y},h_{Z}\}+O(1)$. A standard convention in the field is that all height function equalities and inequalities implicitly account for bounded ambiguity, even if not explicitly stated or if references to $M_{k}$-constants or $O(1)$ terms are omitted.
Particularly, there also exists an intuition of arithmetic complexity behind global function as follows: if a point $P$ is $v$-adically close to $Y$ (though not on $Y$), then after a projective embedding using a very ample multiple of the divisor $A$, the coordinates of $P$ become $v$-adically complex. In reference [Sil87, Proposition 4.2], the author concretely explained this statement. Additionally, this implies that for any arbitrary closed subscheme $Y$, $h_{A}$ dominates $h_{Y}$, as the domain of $h_{A}$ can be extended to contain all of $X(k)$ and there exists a constant $c$ such that
(4) $h_{Y}(P)\leq ch_{A}(P)$
for all $P\in X(k)\setminus Y$.
Another important property of the local height functions ($\lambda_{Y,v})_{v\in M_{k}}$ is that they are bounded from below by an $M_{k}$-constant. This fact immediately shows that the inequality (4) can be formulated for each individual $\lambda_{Y,v}$, i.e.,
(5) $\lambda_{Y,v}(P)\leq ch_{A}(P).$
2.3. Subgeneral position and Seshadri constants
Now we review the definitions of position and Seshadri constants, both in the context of closed subschemes (with analogous definitions applying to divisors).
Definition 2.1.
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$ and let $Y_{1},\ldots,Y_{q}$ be $q$ closed subschemes of $X.$ Suppose $m$ and $\kappa$ are positive integers satisfying $m\geq\dim X\geq\kappa.$
(a). The closed subschemes $\{Y_{1},\ldots,Y_{q}\}$ are taken to be in general position in $X$ if for any subset $I\subset\{1,\ldots,q\}$ with $\sharp I\leq\dim X+1,$
$\operatorname{codim}\left(\bigcap_{i\in I}Y_{i}\cap X\right)\geq\sharp I.$
(b). The closed subschemes $\{Y_{1},\ldots,Y_{q}\}$ are taken to be in $m$-subgeneral position in $X$ if for any subset $I\subset\{1,\ldots,q\}$ with $\sharp I\leq m+1,$
$\dim\left(\bigcap_{i\in I}Y_{i}\cap X\right)\leq m-\sharp I.$
(c). The closed subschemes $\{Y_{1},\ldots,Y_{q}\}$ are in $m$-subgeneral position in $X$ with index $\kappa$ if $\{Y_{1},\ldots,Y_{q}\}$ are in $m$-subgeneral position and for any subset $I\subset\{1,\ldots,q\}$ with $\sharp I\leq\kappa,$
$\operatorname{codim}\left(\bigcap_{i\in I}Y_{i}\cap X\right)\geq\sharp I.$
We now revisit the concept of a Seshadri constant for a closed subscheme relative to a nef Cartier divisor (for additional details, see [HL21, Section 2]).
2.4. Seshadri constants
Let $X$ be a projective variety and $Y$ be a closed subscheme of $X,$ corresponding to a coherent sheaf of ideals $\mathcal{I}_{Y}.$ Consider the sheaf of graded algebras $\mathcal{S}=\oplus_{d\geq 0}\mathcal{I}_{Y}^{d}$, where $\mathcal{I}_{Y}^{d}$ is the $d$-th power of $\mathcal{I}_{Y},$ and $\mathcal{I}_{Y}^{0}=\mathcal{O}_{X}$ by convention. Then $\tilde{X}:=Proj\mathcal{S}$ is called the blowing-up of $X$ with respect to $\mathcal{I}_{Y},$ in other words, the blowing-up of $X$ along $Y.$
Let $\pi:\tilde{X}\to X$ be the blow-up of $X$ along $Y$. It is a fact that the inverse image ideal sheaf ~$\tilde{\mathcal{I}}_{Y}=\pi^{-1}\mathcal{I}_{Y}\cdot\mathcal{O}_{\tilde{X}}$ is an invertible sheaf on $\tilde{X}$.
Now we have the following notion:
Definition 2.2.
Let $Y$ be a closed subscheme of a projective variety $X$ and let $\pi:\tilde{X}\to X$ be the blowing-up of $X$ along $Y$. Let $A$ be a nef Cartier divisor on $X$. The Seshadri constant $\epsilon_{Y}(A)$ of $Y$ with respect to $A$ is the real number defined by
[…]
where $E$ is an effective Cartier divisor on $\tilde{X}$ whose associated invertible sheaf is the dual of $\pi^{-1}\mathcal{I}_{Y}\cdot\mathcal{O}_{\tilde{X}}$.
In addition, we define the vector space $L(D)=\{f\in\overline{k}(X)\mid\text{div}(f)\geq-D\}$, where $\text{div}(f)$ denotes the divisor associated with $f$. The dimension of this vector space over $\overline{k}$ is denoted by $l(D)=\dim_{\overline{k}}L(D)$. We use the symbol $\sim$ to indicate linear equivalence of divisors and $\equiv$ to denote numerical equivalence of divisors.
3. Proof of Theorem 1.7 and Remark 1.8
In this section we disassemble our proof precisely into the following implications (along the lines of [Lev14]):
  • (a): Schmidt’s theorem for numerically equivalent ample divisors
$\Downarrow\mathrm{(i)}$
  • (b): (a) + controlling the dimension of the exceptional set $Z$
$\mathbf{item}$(i): is achieved through repeated applications of (a) to exceptional sets.
$\Downarrow\mathrm{(ii)}$
  • (c): (b) for algebraic points of bounded degree on $X$
$\mathbf{item}$(ii): is exploited by reducing Diophantine approximation problems for algebraic points to those for rational points on symmetric powers of $X$.
In brief, after proving the theorem of type (a), we can derive the results stated in (b) and (c) immediately. As for the proof of type (a), our first method largely present the intensity and logical relation between several different conditions even though more complicated.
First we explore some necessary parameters. It is noteworthy that if not otherwise indicated, in our formula we default to the Seshadri constants $\epsilon_{D_{i,v}}(A)$ corresponding to $d_{i,v}$ successively, rather than its initial definition. Uniform abbreviations and symbol system consistent with [Lev14] and let $c_{i,v}\equiv 1$ in this part:
Given the setup, consider a projective variety $X$ of dimension $n$ defined over a number field $k$, with a finite set of places $S$ of $k$. Let $m\geq n$ be positive integers. For each place $v\in S$, we have $m+1$ effective Cartier divisors $D_{0,v},D_{1,v},\ldots,D_{m,v}$ on $X$ that are in $m$-subgeneral position, i.e., $\cap_{i=0}^{m}\operatorname{Supp}D_{i,v}=\varnothing$. We define a divisor $D$ as the sum of all these divisors over all places in $S$, i.e., $D=\sum_{v\in S}\sum_{i=0}^{m}D_{i,v}$. Additionally, we are given an ample Cartier divisor $A$ on $X$ and positive integers $d_{i,v}$ such that each divisor $D_{i,v}$ is numerically equivalent to $d_{i,v}A$, i.e., $D_{i,v}\equiv d_{i,v}A$ for all $i$ and $v\in S$ ( $\epsilon_{D_{i,v}}(A)\Leftrightarrow\gamma_{D_{i,v}}=\frac{1}{d_{i,v}}$, as seen in Part I, Perspective-ii). For $c\in\mathbb{R}$ and $\delta\in\mathbb{N}$, define the geometric exceptional set $Z=Z\left(c,\delta,m,n,k,S,X,A,\{D_{i,v}\}\right)$ to be the smallest Zariski-closed subset of $X$ such that the set
(6) $\left\{P\in X(\overline{k})\setminus\operatorname{Supp}D\mid\sum_{v\in S}\sum_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\sum_{i=0}^{m}\epsilon_{D_{i,v}}(A)\lambda_{D_{i,v},w}(P)>ch_{A}(P)\right\}\setminus Z$
satisfied for $[k(P):k]\leq\delta$, is finite for all choices of the Weil and height functions.
Now define
$C(\delta,l,m,n)=\sup\{c\mid\dim Z\left(c,\delta,m,n,k,S,X,A,\{D_{i,v}\}\right)\geq l\\ \text{ for some $k$, $S$, $X$, $A$, $\{D_{i,v}\}$ as above}\}.$
Furthermore, it would be beneficial to define and examine specific instances concerning rational points for it represents the initial factor without $\delta$:
$A(m,n)=C(1,n,m,n),B(l,m,n)=C(1,l,m,n).$
In order to prove Theorem 1.7, we require the following inequalities due to [HL25] (The proof process is consistent so we will not interpret further here), which is a generalization of Chebyshev’s inequality (see also work of Jensen [MPF93]).
Lemma 3.1.
Let Weil functions family $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}\geq 0$, set $\lambda_{i}=\log\boxdot_{i}$ and $b_{1},\ldots,b_{n},c_{1},\ldots,c_{n}$ be nonnegative real numbers as sets of indicators. Let $i_{0}$ be the smallest one of the index $i$ such that $c_{i}\not=0$ (assume that there exists at least one satisfied), then
(7) $\log\prod_{i=1}^{n}\boxdot^{b_{i}}_{i}\geq\log\prod_{i=1}^{n}\boxdot^{\left(\min_{i_{0}\leq j\leq n}\frac{\sum_{i=1}^{j}b_{i}}{\sum_{i=1}^{j}c_{i}}\right)c_{i}}_{i}.$
Corollary 3.2.
Let Weil functions family $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}\geq 0$, set $\lambda_{i}=\log\boxdot_{i}$ and $b_{1},\ldots,b_{n},c_{1},\ldots,c_{n}$ be nonnegative real numbers as sets of indicators. Assume that $b_{1}\not=0$. Then
$\log\prod_{i=1}^{n}\boxdot^{\left(\max_{1\leq j\leq n}\frac{\sum_{i=1}^{j}c_{i}}{\sum_{i=1}^{j}b_{i}}\right)b_{i}}_{i}\geq\log\prod_{i=1}^{n}\boxdot^{c_{i}}_{i}.$
[…]
$\epsilon_{D_{i,v}}(A)\lambda_{D_{i,v},v}(P)=\log\boxdot^{\epsilon_{D_{i,v}}(A)}_{D_{i,v},v}(P).$
We now conduct the proof of Theorem 1.7 (invoking the same method in [HL25] to attain the result of effective Cartier divisors) and Remark 1.8.
Proof.
Part I: Schmidt’s theorem for numerically equivalent ample divisors.
$\mathbf{Perspective}-\mathrm{i}$: Let $v\in S$, $c_{i,v}$ $(1\leq i\leq q)$ be nonnegative real numbers and $D_{i,v}$ be effective Cartier divisors. Set $q\geq m+1$ in this part.
(i) Normalized in a two-step process (In view of the Seshadri constants (original definition) $\epsilon_{D_{1,v}}(A),\ldots,\epsilon_{D_{q,v}}(A)$ may be irrelevant distinct positive real numbers):
(a) Rationalization: firstly for every $i=1\ldots,q$, let $\tau_{1,v},\ldots,\tau_{q,v}$ be real numbers in the interval $[0,1)$ so that
$(1+\tau_{1,v})\epsilon_{D_{1,v}}(A),\ldots,(1+\tau_{q,v})\epsilon_{D_{q,v}}(A)$
are rational numbers. If $\epsilon_{D_{i,v}}(A)$ is already a rational number then choose $\tau_{i,v}=0$.
(b) Integer conversion: For a sufficiently divisible integer $c_{v}$ (independent of $i$) such that
$c_{v}(1+\tau_{1,v})\epsilon_{D_{1,v}}(A),\ldots,c_{v}(1+\tau_{q,v})\epsilon_{D_{q,v}}(A)$
are all integers.
By means of the multiplication of the effective Cartier divisors $D_{i,v}$, we set
$\tilde{D}_{i,v}:=c_{v}(1+\tau_{i,v})\epsilon_{D_{i,v}}(A)D_{i,v},$
we have that
(8) $\epsilon_{\tilde{D}_{i,v}}(A)=\frac{1}{c_{v}(1+\tau_{i,v})}$
for $i=1,\ldots,q$, and therefore
$\frac{1}{c_{v}}\leq 2\epsilon_{\tilde{D}_{i,v}}(A)\Rightarrow\frac{1}{c_{v}}\leq 2\min_{i}\epsilon_{\tilde{D}_{i,v}}(A).$
For each Seshadri constant and for all $i,i^{\prime}=1,\ldots,q$, we have
$\ |\epsilon_{\tilde{D}_{i,v}}(A)-\epsilon_{\tilde{D}_{i^{\prime},v}}(A)|$
$=$ $\ \left|\frac{1}{c_{v}(1+\tau_{i,v})}-\frac{1}{c_{v}(1+\tau_{i^{\prime},v})}\right|$
$=$ $\ \frac{1}{c_{v}}\left|\frac{1}{(1+\tau_{i,v})}-\frac{1}{(1+\tau_{i^{\prime},v})}\right|$
$=$ $\ \frac{1}{c_{v}}\left|\frac{\tau_{i^{\prime},v}-\tau_{i,v}}{(1+\tau_{i^{\prime},v})(1+\tau_{i,v})}\right|$
$\leq$ $\ \max_{k}\frac{\tau_{k,v}}{c_{v}}$
(9) $\leq$ $\ \max_{k}\tau_{k,v}\cdot 2\min_{l}\epsilon_{\tilde{D}_{l,v}}(A)$
(ii) Intersections of the normalized divisors: set
$\tilde{D}_{I,v}=\bigcap_{i\in I}\tilde{D}_{i,v}.$
In fact, $\tilde{D}_{I,v}$ here are defined to be closed subschemes by using the Lemma 2.2 in [Sil87]. We simply refer to them as divisors in the sequel. Fix a point $P$ outside of the support of the given divisors. Let $\{i_{1,v},\ldots,i_{q,v}\}=\{1,\ldots,q\}$ be such that:
(10) $\log\boxdot_{\tilde{D}_{i_{1,v},v},v}(P)\geq\ldots\geq\cdots\geq\log\boxdot_{\tilde{D}_{i_{q,v},v},v}(P)\geq 0.$
We set $I_{j,v}=\{i_{1,v},\ldots,i_{j,v}\}$. Let $l_{v}$ be the largest index such that
$\cap_{j=1}^{l_{v}}\tilde{D}_{i_{j,v},v}\neq\varnothing.$
With the property (2) and the sequence of inequalities (10) we have
$\log\boxdot_{\tilde{D}_{i_{j,v}v},v}(P)=\min_{j^{\prime}\leq j}\log\boxdot_{\tilde{D}_{i_{j^{\prime},v},v},v}(P)=\log\boxdot_{\tilde{D}_{I_{j,v},v},v}(P).$
In particular, for $j>l_{v}$, $\log\boxdot_{\tilde{D}_{i_{j,v},v},v}(P)$ is bounded by a constant independent of $P$, and
$\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot^{c_{i,v}\epsilon_{\tilde{D}_{i,v}}(A)}_{\tilde{D}_{i,v},v}(P)\leq\log\prod_{v\in S}\prod_{j=1}^{l_{v}}\boxdot^{c_{i_{j},v}\epsilon_{\tilde{D}_{i_{j,v},v}}(A)}_{\tilde{D}_{I_{j,v},v},v}(P).$
Let $b_{j,v}=\operatorname{codim}\tilde{D}_{I_{j,v},v}$ and set $b_{0,v}=0$. Set $\epsilon_{v}(A)=\max_{i}\epsilon_{\tilde{D}_{i,v}}(A)$. Apply Corollary 3.2 in this condition with
$\lambda_{1}$ $=\log\boxdot^{\epsilon_{v}(A)}_{\tilde{D}_{I_{1,v},v},v}(P)\geq\cdots\geq\lambda_{l_{v}}=\log\boxdot^{\epsilon_{v}(A)}_{\tilde{D}_{I_{l_{v},v},v},v}(P)\geq 0,$
$b_{1}$ $=b_{1,v}-b_{0,v},\ldots,b_{l_{v}}=b_{l_{v},v}-b_{l_{v}-1,v},$
$c_{j}$ $=c_{i_{j},v}.$
Using the telescoping property for the $b_{i}$,
(11) $\log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{\left(\max_{j}\frac{\sum_{j^{\prime}=1}^{j}c_{i_{j^{\prime}},v}}{b_{j,v}}\right)(b_{s,v}-b_{s-1,v})\epsilon_{v}(A)}_{\tilde{D}_{I_{s,v},v},v}(P)\geq\log\prod_{v\in S}\prod_{j=1}^{l_{v}}\boxdot^{c_{i_{j},v}\epsilon_{v}(A)}_{\tilde{D}_{I_{j,v},v},v}(P).$
According to [Laz04, Example 5.4.11] (This is expressed in terms of the $s$invariant, which is the reciprocal of the Seshadri constant), the Seshadri constant of an intersection is bounded below by the minimum of the Seshadri constants of the factors of the intersection. Therefore, for any $j\in\{1,\ldots,l_{v}\}$, combined with the (9), we have the inequality
(12) $\epsilon_{\tilde{D}_{I_{j,v},v}}(A)\geq\min\{\epsilon_{\tilde{D}_{i_{1},v}}(A),\ldots,\epsilon_{\tilde{D}_{i_{j},v}}(A)\}\geq\epsilon_{v}(A)-\max_{k}\tau_{k,v}\cdot 2\min_{l}\epsilon_{\tilde{D}_{l,v}}(A),$
Then (11) implies
$\log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{\left(\max_{j}\frac{\sum_{j^{\prime}=1}^{j}c_{i_{j^{\prime}},v}}{b_{j,v}}\right)(b_{s,v}-b_{s-1,v})\left(\epsilon_{\tilde{D}_{I_{s,v},v},v}(A)+\max_{k}\tau_{k,v}\cdot 2\min_{l}\epsilon_{\tilde{D}_{l,v}}(A)\right)}_{\tilde{D}_{I_{s,v},v},v}(P)$
$\geq$ $\log\prod_{v\in S}\prod_{j=1}^{l_{v}}\boxdot^{c_{i_{j},v}\epsilon_{v}(A)}_{\tilde{D}_{I_{j,v},v},v}(P)\geq\log\prod_{v\in S}\prod_{j=1}^{l_{v}}\boxdot^{c_{i_{j},v}\epsilon_{\tilde{D}_{i_{j,v},v}}(A)}_{\tilde{D}_{I_{j,v},v},v}(P).$
If we construct a new list by repeating the divisors $\tilde{D}_{I_{s,v},v}$ exactly $b_{s,v}-b_{s-1,v}$ times (and omitting them if $b_{s,v}-b_{s-1,v}=0$), then the resulting divisors will be in general position. By Theorem 1.1, there exists a proper Zariski closed subset $Z$ of $X$ such that
$\log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{(b_{s,v}-b_{s-1,v})\epsilon_{\tilde{D}_{I_{s,v},v}}(A)}_{\tilde{D}_{I_{s,v},v},v}(P)<(n+1+\epsilon^{\prime})h_{A}(P)$
for all points $P\in X(k)\setminus Z$. We also find that by (8) and additivity for local heights, for $i=1,\ldots,q$, we have
$\log\boxdot^{\epsilon_{\tilde{D}_{i,v}}(A)}_{\tilde{D}_{i,v},v}(P)=\log\boxdot^{\frac{1}{c_{v}(1+\tau_{i,v})}\cdot c_{v}(1+\tau_{i,v})\epsilon_{D_{i,v}}(A)}_{D_{i,v},v}(P)=\log\boxdot^{\epsilon_{D_{i,v}}(A)}_{D_{i,v},v}(P).$
For a given $\epsilon^{\prime\prime}>0$, by selecting the quantities $\tau_{k,v}$ sufficiently small (with the choice depending only on data involving the $D_{i,v}$, including the finite number of pertaining domination inequalities of the form (5) for the $D_{i,v}$), we get
$\log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{(b_{s,v}-b_{s-1,v})(\max_{k}\tau_{k,v}\cdot 2\min_{l}\epsilon_{\tilde{D}_{l,v}}(A))}_{\tilde{D}_{I_{s,v},v},v}(P)$
$\leq$ $\ \log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{(b_{s,v}-b_{s-1,v})(\max_{k}\tau_{k,v}\cdot 2\epsilon_{\tilde{D}_{i_{s,v},v}}(A))}_{\tilde{D}_{i_{s,v},v},v}(P)$
$\leq$ $\ \log\prod_{v\in S}\prod_{s=1}^{l_{v}}\boxdot^{(b_{s,v}-b_{s-1,v})(\max_{k}\tau_{k,v}\cdot 2\epsilon_{D_{i_{s,v},v}}(A))}_{D_{i_{s,v},v},v}(P)$
$<$ $\ \epsilon^{\prime\prime}h_{A}(P).$
Taking $W=\operatorname{Supp}\tilde{D}_{I_{j,v},v}\neq\varnothing$, we have
$\frac{\sum_{j^{\prime}=1}^{j}c_{i_{j^{\prime}},v}}{b_{j,v}}\leq\frac{\alpha_{v}(W)}{\operatorname{codim}W}.$
It should be noted that there are only finitely many possible combinations for indices $i_{1,v},\ldots,i_{l_{v},v}$. Consequently, only finitely many applications of Theorem 1.1 are required. Taking all of the above into consideration, there exists a proper Zariski closed set $Z$ of $X$ such that
(13) $\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot^{c_{i,v}\epsilon_{D_{i,v}}(A)}_{D_{i,v},v}(P)\leq\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{A}(P)$
for all points $P\in X(k)\setminus Z$, as desired.
The core of the next proof section (basically follows [Lev14]) is iteratively applying the refined version of (13). Now, it shifts to prove (13) holds for $D_{i,v}\sim d_{i,v}A$:
Let $N$ be a positive integer satisfying $NA$ is very ample and $d_{i,v}$ divides $N$ for all $i$ and $v\in S$. Define $M=l(NA)-1$. Let $\phi:X\to\mathbb{P}^{M}$ be the embedding of $X$ in $\mathbb{P}^{M}$ associated to $NA$. Since $\frac{N}{d_{i,v}}D_{i,v}\sim NA$, there exists some hyperplanes $H_{i,v}\in\mathbb{P}^{M}$ defined over $k$ such that $\frac{N}{d_{i,v}}D_{i,v}=\phi^{*}H_{i,v}$. For $P\in X(k)\setminus\operatorname{Supp}D_{i,v}$, by functoriality and additivity of Weil functions, we have
$\log\boxdot_{H_{i,v},v}(\phi(P))=\log\boxdot_{\frac{N}{d_{i,v}}D_{i,v},v}(P)=N\log\boxdot^{\frac{1}{d_{i,v}}}_{D_{i,v},v}(P).$
We also find $h(\phi(P))=Nh_{A}(P)+O(1)$ for $P\in X(k)$. Plugging these identities into (13) in the case of hyperplanes with $c_{i,v}=1$, we can get the desired result.
Ultimately, we demonstrate that the statement remains valid even if we replace linear equivalence with numerical equivalence. This conclusion is supported by the following result due to Matsusaka [Mat58] in algebraic geometry:
Lemma 3.3 (Matsusaka).
Let $A$ be an ample Cartier divisor on a projective variety $X$. For any Cartier divisor $D$, there exists a positive integer $N_{0}$ satisfying for all $N\geq N_{0}$ such that $D\equiv NA$, $D$ is very ample.
The key point is to form the following family of divisors $Q_{i,v}$ to replace $D_{i,v}$:
It is sufficient to prove the case where $d_{i,v}=1$ for all $i$ and all $v$ by setting the least common multiple of the integers $d_{i,v}$, i.e., $D_{i,v}\equiv A$ for all $i$ and all $v$. Let $N_{0}$ be the integer in lemma 3.3 for $A$. By standard properties of Weil functions and heights we have mentioned before, there exists an integer $N>N_{0}$ such that
(14) $\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot^{\frac{N_{0}}{N}}_{D_{i,v},v}(P)<\epsilon h_{A}(P)+O(1)$
for all $P\in X(k)\setminus\cup_{i,v}\operatorname{Supp}D_{i,v}$. By the selection of $N_{0}$, it follows that $NA-(N-N_{0})D_{i,v}$ is very ample for all $i$ and $v\in S$. According to this fact, there exist effective divisors $E_{i,v}$ such that $NA\sim(N-N_{0})D_{i,v}+E_{i,v}$, and for all $v\in S$, the divisors $(N-N_{0})D_{1,v}+E_{1,v},\ldots,(N-N_{0})D_{q,v}+E_{q,v}$ and $D_{1,v},\ldots,D_{q,v}$ have the same characteristic of position. We now apply (13) with $c_{i,v}=1$ to the linearly equivalent divisors $Q_{i,v}=(N-N_{0})D_{i,v}+E_{i,v}$, $i=1,\ldots,q$, $v\in S$, along with $NA$.
$\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot_{(N-N_{0})D_{i,v}+E_{i,v},v}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{NA}(P)$
for all $P\in X(k)\setminus Z$ for some proper Zariski-closed subset $Z$ of $X$ containing the supports of all $E_{i,v}$, $v\in S$, $i=1,\ldots,q$, is established. Since $\log\boxdot_{E_{i,v},v}$ is bounded from below outside of the support of $E_{i,v}$, we obtain
$\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot^{\left(1-\frac{N_{0}}{N}\right)}_{D_{i,v},v}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{A}(P)+O(1)$
for all $P\in X(k)\setminus Z$. Inserting (14) into above formula we deduce that
(15) $\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot_{D_{i,v},v}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+2\epsilon\right)h_{A}(P)+O(1)$
for all $P\in X(k)\setminus Z$. Given that there are only finitely many $k$-rational points with bounded height (with respect to $A$), we can ensure that equation (15) holds without the $O(1)$ term by adding a finite number of points to $Z$.
$\mathbf{Perspective}-\mathrm{ii}$: Another proof version is by using an immediate corollary of Theorem 1.4 by simplifying the Seshadri constant via its inherent properties, i.e.
[…]
when $D$ is an effective Cartier divisor, thus $\tilde{X}=X$. Thereby we have,
Corollary 3.4.
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $D_{1,v},\ldots,D_{q,v}$ be effective Cartier divisors of $X$, defined over $k$, set $\gamma_{D_{i,v}}$ be rational numbers such that $A-\gamma_{D_{i,v}}D_{i,v}$ is a nef $\mathbb{Q}$-divisor for all $i$ and for all $v\in S$, $A$ is an ample Cartier divisor, then there exist a proper Zariski-closed subset $Z\subset X$ such
$\log\prod_{v\in S}\prod_{i=1}^{q}\boxdot^{c_{i,v}\gamma_{D_{i,v}}}_{D_{i,v},v}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{A}(P)$
for all points $P\in X(k)\setminus Z$ and $\epsilon>0$, is established with $c_{i,v}\in\mathbb{R}^{+}\cup\{0\}$.
It can plainly set $\epsilon_{D_{i,v}}(A)\Leftrightarrow\gamma_{D_{i,v}}=\frac{1}{d_{i,v}}$ (i.e.-$\mathbf{corresponding}$) to yield (15).
Part II: An inequality within parameters and controlling the dimension of $Z$.
In consideration of the numerically equivalent ample divisors, we will henceforth assume that $c_{i,v}=1$ for all $i$ and $v$. Now it is direct to verify that (i.e. the proof of the relatively complex side of the index in the inequality of [Lev14])
$C(\delta,l,m,n)\leq B(l,\delta m,\delta n)\leq\max_{l\leq j\leq\delta n}A(\delta m,j).$
[…]
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot^{\epsilon_{D_{{i,v}|W}}(A)}_{D_{{i,v}|W},w}(P)>(c-\epsilon)h_{A|_{W}}(P)$
is Zariski-dense in $W$. It shows that $A(m,j)\geq c$ which implies the right-hand side.
To simplify the exposition and without loss of generality, we will replacing $\left(\prod_{j\neq i}\epsilon_{D_{i,v}}(A)\right)D_{i,v}$ by $D_{i,v}$ and $\left(\prod_{j=1}^{q}\epsilon_{D_{i,v}}(A)\right)A$ by $A$, thus using $\epsilon_{D_{i,v}}(A)=1$ as a prerequisite. Forming the set $Z$ defined to be a Zariski-closure of the set
$\left\{P\in X(\overline{k})\setminus\operatorname{Supp}D\mid\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D_{i,v},w}(P)>(B(l,\delta m,\delta n)+\epsilon)h_{A}(P)\right\}$
[…]
(i) Notation and Definitions: Let $X^{\delta}$ denote the $\delta$th power of $X$, which is the Cartesian product of $X$ with itself $\delta$ times. Let $X^{(\delta)}$ denote the $\delta$-th symmetric power of $X$, defined as the quotient $X^{\delta}/S_{\delta}$, where $S_{\delta}$ is the symmetric group on $\delta$ elements acting by permuting the coordinates of $X^{\delta}$. Let $\phi:X^{\delta}\to X^{(\delta)}$ be the natural projection map that takes a point in $X^{\delta}$ to its equivalence class in $X^{(\delta)}$.
(ii) Points and Conjugates: For a point $P\in X(\overline{k})$ satisfying $[k(P):k]=\delta$, let $P=P_{1},P_{2},\ldots,P_{\delta}$ denote the $\delta$ distinct conjugates of $P$ over $k$. Define the map $\psi:X(\overline{k})\to X^{\delta}$ by $\psi(P)=(P_{1},P_{2},\ldots,P_{\delta})$. We can also verify that the point $\phi(\psi(P))\in X^{(\delta)}$ is $k$-rational.
(iii) Push-forward and Pullback: Let $\pi_{i}:X^{\delta}\to X$ denote the $i$th projection map, which projects a point in $X^{\delta}$ onto its $i$-th coordinate. Define $D^{*}_{i,v}=\phi_{*}\pi_{1}^{*}(D_{i,v})$ and $D^{*}=\sum_{v\in S}\sum_{i=1}^{q}D^{*}_{i,v}$. Also define $A^{*}=\phi_{*}\pi_{1}^{*}(A)$.
(iv) Properties of Divisors: Note that $\phi^{*}\phi_{*}\pi_{1}^{*}D_{i,v}=\sum_{j=1}^{\delta}\pi_{j}^{*}D_{i,v}$. Since the $D_{i,v}$ are in $m$-subgeneral position, the $D^{*}_{i,v}$ are in $m\delta$-subgeneral position. It can be verified that $A^{*}$ is also ample. Consequently, $D^{*}_{i,v}$ is effective and ample for all $i$ and $v$. If $D_{i,v}\equiv D_{j,v}\equiv A$ for all $i$ and $j$ , then $D^{*}_{i,v}\equiv D^{*}_{j,v}\equiv A^{*}$ either.
Thus by the definition of $B(l,\delta m,\delta n)$, we have the inequality
(16) $\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D^{*}_{i,v},w}(P)<(B(l,\delta m,\delta n)+\epsilon)h_{A^{*}}(P)$
$\quad\forall P\in X^{(\delta)}(k)\setminus(Z^{\prime}\cup\operatorname{Supp}D^{*}),$
for some Zariski-closed subset $Z^{\prime}$ of $X^{(\delta)}$ satisfying $\dim Z^{\prime}\leq l-1$.
For $P\in X(\overline{k})\setminus\operatorname{Supp}D$, $[k(P):k]=\delta$, we also have, up to $O(1)$,
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D^{*}_{i,v},w}(\phi(\psi(P)))$
[…]
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{\phi_{*}\pi_{1}^{*}D_{i,v},w}(\phi(\psi(P)))$
[…]
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{\phi^{*}\phi_{*}\pi_{1}^{*}D_{i,v},w}(\psi(P))=\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{\sum_{j=1}^{\delta}\pi_{j}^{*}D_{i,v},w}(\psi(P))$
[…]
$\sum_{j=1}^{\delta}\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{\pi_{j}^{*}D_{i,v},w}(\psi(P))$
[…]
$\sum_{j=1}^{\delta}\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D_{i,v},w}(\pi_{j}\psi(P))=\sum_{j=1}^{\delta}\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D_{i,v},w}(P_{j})$
[…]
$\delta\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D_{i,v},w}(P).$
Identically, we also have $h_{A}^{*}(\phi(\psi(P)))=\delta h_{A}(P)+O(1)$. Consider $Z^{*}\subset Z(\overline{k})$ as a Zariski-dense set of points in $Z$ satisfying $[k(P):k]=\delta$. It follows that for an appropriate choice of the functions,
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D^{*}_{i,v},w}(\phi(\psi(P)))>(B(l,\delta m,\delta n)+\epsilon)h_{A^{*}}(\phi(\psi(P)))$
for all $P\in Z^{*}$. From (16), we have $\dim\phi(\psi(Z^{*}))\leq l-1$. Since $\phi$ is a finite map and $\pi_{1}(\psi(Z^{*}))=Z^{*}$, $\dim\phi(\psi(Z^{*}))\geq\dim Z^{*}=\dim Z$. Thus $\dim Z\leq l-1$, as desired. It shows that $C(\delta,l,m,n)\leq B(l,\delta m,\delta n)$. Based on the above construction and discussion, we can also obtain $B(l,\delta m,\delta n)\leq\max_{l\leq j\leq\delta n}A(\delta m,j)$.
Remark 3.5.
The purpose of proving the inequality is twofold: Fill the void that this theorem lack of conforms to the form of our results and concretely reveal its intrinsic characteristic and application performance to refine our presentation in the next part.
Part III: Part II for algebraic points of bounded degree.
Leveraging the result in Part I and integrate with the Part II, we can obtain Theorem 1.7 with the factor of $a_{\delta}+\epsilon$ immediately (along the same lines, even if $c_{i,v}\not\equiv 1$, the formula still holds). Now we adjust the factor $a_{\delta}$ on the right side to $a_{\Delta,\delta}$, where $a_{\Delta,\delta}$ here is defined by
$\left[(n+1)\max_{v\in S}\delta_{\mathcal{D}_{v}}(i.e.-\Delta)\right]_{w.r.t-\delta}$
while
$\delta_{\mathcal{D}_{v}}=\max_{I\subset\{1,\ldots,q\}}\max\left\{1,\frac{\#I}{\operatorname{codim}\cap_{i\in I}\operatorname{Supp}D_{i,v}}\right\}=\max\left\{1,\max_{\varnothing\subsetneq W\subsetneq X}\frac{\alpha_{v}(W)}{\operatorname{codim}W}\right\},$
where we set $\operatorname{codim}\varnothing=\infty$.
To explain clearly, we are here to show that above two definitions are equivalent:
Assume that $\tilde{W}$ is a subvariety of $X$ where the ratio $\frac{\alpha_{v}(W)}{\operatorname{codim}W}$ achieves its maximum value precisely when $W=\tilde{W}$. By reordering if necessary, we can assume that $\tilde{W}\subset D_{i,v}$ for $i=1,\dots,\alpha_{v}(\tilde{W})$. Define $\bar{W}=\bigcap_{i=1}^{\alpha_{v}(\tilde{W})}D_{i,v}$. It is evident that $\tilde{W}\subset\bar{W}$, which implies that $\operatorname{codim}\tilde{W}\geq\operatorname{codim}\bar{W}$. Additionally, since $\tilde{W}\nsubseteq D_{i,v}$ for all $i>\alpha_{v}(\tilde{W})$, it follows that $\bar{W}\nsubseteq D_{i,v}$ for all $i>\alpha_{v}(\tilde{W})$. Therefore, we conclude that $\alpha_{v}(\tilde{W})=\alpha_{v}(\bar{W})$. Thus, we deduce that
$\frac{\alpha_{v}(\bar{W})}{\operatorname{codim}\bar{W}}\geq\frac{\alpha_{v}(\tilde{W})}{\operatorname{codim}\tilde{W}}.$
By our assumption for $\tilde{W}$, we get
$\frac{\alpha_{v}(\bar{W})}{\operatorname{codim}\bar{W}}=\frac{\alpha_{v}(\tilde{W})}{\operatorname{codim}\tilde{W}}.$
Our claim follows from this observation which means we only need to consider those $W$ that are the intersections of some $D_{i,v}$.
In conjunction with the inequality and assumptions in Part II we have,
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $D_{1,v},\ldots,D_{q,v}$ $(q=m+1)$ be effective (ample) Cartier divisors of $X$, defined over $k$, set Seshadri constants $\epsilon_{D_{i,v}}(A)$ corresponding to positive integers $d_{i,v}$ (i.e.-$\Leftrightarrow\gamma_{D_{i,v}}=\frac{1}{d_{i,v}}$) such that $D_{i,v}\equiv d_{i,v}A$ for all $i$ and for all $v\in S$, while $A$ is an ample Cartier divisor, then
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot^{\epsilon_{D_{i,v}}(A)}_{D_{i,v},w}(P)<\left(a_{\Delta,\delta}+\epsilon\right)h_{A}(P)+O(1)$
for all points $P\in X(\bar{k})\setminus\cup_{i,v}\operatorname{Supp}D_{i,v}$ and $\epsilon>0$, as desired.
At this point, we have proven the Remark 1.8 in general form. In comparison with theorem 1.7, due to the particularity of other cases, we need a unified, intuitive and more general representation, that is the rationale behind we adopt the form of our following main result for resolution:
Using the definition if $D_{i,v}$ are taken to be effective Cartier divisors in $m$-subgeneral position, then
$\Delta=\max_{v\in S}\delta_{\mathcal{D}_{v}}\leq m-n+1.$
The inequality in Part II shows that,
$C(\delta,1,m,n)\leq B(1,\delta m,\delta n)\leq\max_{1\leq j\leq\delta n}A(\delta m,j)\leq\max_{1\leq j\leq\delta n}\left(\delta m-j+1\right)(j+1):a_{\Delta,\delta}$
Consider $\left(\delta m-j+1\right)(j+1)$ as a quasi-quadratic function with the value of variable be integers, based on the most basic knowledge, our results split into two cases:
  • (i)For $m>2n$,
    $a_{\Delta,\delta}=n({\delta^{2}}m-{\delta^{2}}n)+{\delta}m+1$
  • (ii)For $n\leq m\leq 2n$, if $j=\left[\frac{\delta m}{2}\right]_{Int}\Leftrightarrow i$, we can obtain the maximum:
    $a_{\Delta,\delta}=-i^{2}+\delta m\cdot i+{\delta}m+1$
And if $D_{i,v}$ are divisors in $m$-subgeneral position with index $\kappa$, then
$\Delta\leq\frac{m-n}{\kappa}+1,$
which implies that,
$C(\delta,1,m,n)\leq B(1,\delta m,\delta n)\leq\max_{1\leq j\leq\delta n}A(\delta m,j)\leq\max_{1\leq j\leq\delta n}\left(\frac{\delta m-j}{\kappa}+1\right)(j+1)$
The same operation as above, we affirmative that the factors below established:
  • (i)For $m\geq\lceil 2n+\frac{1-\kappa}{\delta}\rceil$,
    $a_{\Delta,\delta}=\frac{n}{\kappa}\left({\delta^{2}}m-{\delta^{2}}n+\delta(\kappa-1)\right)+\frac{{\delta}m}{\kappa}+1$
  • (ii)For $n\leq m\leq\lfloor 2n+\frac{1-\kappa}{\delta}\rfloor$, $j=\left[\frac{\delta m+(\kappa-1)}{2}\right]_{Int}\Leftrightarrow i$ in this situation:
    $a_{\Delta,\delta}=-\frac{i^{2}}{\kappa}+\left(\frac{\delta m+(\kappa-1)}{\kappa}\right)i+\frac{{\delta}m}{\kappa}+1$
thus we completely end our proof as promised by Remark 1.8. ∎
As a powerful concept to measure positivity, Seshadri constants release the tight of restricted conditions in a way, this create a version for us to loosen the assumption of our results, we can get the following corollary by modifying some points:
Corollary 3.6.
Let $X$ be a projective variety of dimension $n$ defined over a number field $k$, and let $S$ be a finite set of places of $k$. For each $v\in S$, let $D_{1,v},\ldots,D_{q,v}$ ($q=m+1$) be effective Cartier divisors of $X$ in $m$-subgeneral position, defined over $k$, set $\gamma_{D_{i,v}}$ be rational numbers such that $A-\gamma_{D_{i,v}}D_{i,v}$ is a nef $\mathbb{Q}$-divisor for all $i$ and for all $v\in S$, while $A$ is an ample Cartier divisor, then there exist a proper Zariski-closed subset $Z\subset X$ and $\epsilon>0$ such that
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot^{c_{i,v}\gamma_{D_{i,v}}}_{D_{i,v},w}(P)<\left(a_{\delta}+\epsilon\right)h_{A}(P)+O(1)$
for all points $P\in X(\bar{k})\setminus\cup_{i,v}\operatorname{Supp}D_{i,v}$ satisfying $[k(P):k]\leq\delta$, as desired.
$c_{1,v},\ldots,c_{q,v}$ here are nonnegative real numbers defined in Theorem 1.4. And set
$\alpha_{v}(W)=\sum_{\begin{subarray}{c}i\\ W\subset\operatorname{Supp}D_{i,v}\end{subarray}}c_{i,v}.$
in the definition of $a_{\delta}$, where $W\subset X$ is a closed subset.
Proof.
Let $\gamma_{D_{i,v}}$ be rational numbers such that $A-\gamma_{D_{i,v}}D_{i,v}$ is a nef $\mathbb{Q}$-divisor for all $i$ and for all $v\in S$.
Consider the quantities and divisors $c,l,m,n,k,S,X,A,D_{1,v},\ldots,D_{q,v}$ (where $q=m+1$ and in $m$-subgeneral position), and $D$, as previously defined. For $\delta\in\mathbb{N}$, define the geometric exceptional set $Z=Z\left(c,\delta,m,n,k,S,X,A,\{D_{i,v}\}\right)$ to be the smallest Zariski-closed subset of $X$ such that the set
$\left\{P\in X(\overline{k})\setminus\operatorname{Supp}D\mid\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot_{D_{i,v},w}^{c_{i,v}\gamma_{D_{i,v}}}(P)>ch_{A}(P)\right\}\setminus Z$
satisfied for $[k(P):k]\leq\delta$, is finite for all choices of the Weil and height functions. Analogous definitions as previous section for $C(\delta,l,m,n)$, $B(l,\delta m,\delta n)$ and $A(\delta m,j)$. If for rational points with $\delta=1$, identical discussion and hypothesis, we have that
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot^{c_{i,v}\gamma_{D_{{i,v}|W}}}_{D_{{i,v}|W},w}(P)>(c-\epsilon)h_{A|_{W}}(P)$
is Zariski-dense in $W$. It implies the right-hand side of the inequality in Part II for rational points with $\delta=1$ (this also hold true for algebraic points with $\delta\in\mathbb{N}$).
Let $\epsilon>0$, construct the set $Z$ defined to be a Zariski-closure of the set
$\{P\in X(\overline{k})\setminus\operatorname{Supp}D\mid[k(P):k]=\delta,\\ \log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot^{c_{i,v}\gamma_{D_{i,v}}}_{D_{i,v},w}(P)>(B(l,\delta m,\delta n)+\epsilon)h_{A}(P)\}$
To simplify the proof and without loss of generality, replacing $\left(\prod_{j\neq i}\gamma_{D_{i,v}}\right)D_{i,v}$ by $D_{i,v}$ and $\left(\prod_{j=1}^{q}\gamma_{D_{i,v}}\right)A$ by $A$, i.e.-$\gamma_{D_{i,v}}=1$ as usual.
With the same mapping construction as in Part II, i.e. $D^{*}_{i,v}=\phi_{*}\pi_{1}^{*}(D_{i,v})$ and $A^{*}=\phi_{*}\pi_{1}^{*}(A)$. The key issue is to prove that such mapping is flat, in order to obtain that $A^{*}-\gamma_{D_{i,v}}D_{i,v}^{*}$ is $\mathbb{Q}$-nef divisor either. Under our assumption, the map $\phi$ is a finite quotient map, which is flat with respect to the action of the symmetric group $S_{\delta}$ via its free action on $X^{\delta}$. Thus the push-forward map $\phi_{*}$ preserves flatness. Also, map $\pi_{i}$ is a standard flat projection map in algebraic geometry. Thereby we derive our conclusion. The residual procedure adhere to Part II. In conjunction with the Corollary 3.4 and $a_{\delta}$ (also apply to $a_{\Delta,\delta}$ when $c_{i,v}\equiv 1$), we have our result. ∎
  • @itemi@itemi@itemi@itemi
    Note that our result could be easily extended to the formula:
    $\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\max_{I}\prod_{i\in I}\boxdot^{c_{i}\gamma_{D_{i}}}_{D_{i,w}}(P)<\left(a_{\delta}+\epsilon\right)h_{A}(P)+O(1)$
by invoking the Theorem 2.5.8 in [Voj87] as an equivalent alternative, where $I$ loops through all subsets of $\left\{1,\dots,q\right\}$ ($q\geq m+1$) such that $\left\{D_{i}\right\}$, $i\in I$ are in $m$-subgeneral position. The following proof is aligned with Theorem 5.1, [Lev14].
Proof.
Since the divisors $D_{i}$ are in $m$-subgeneral position, any point $P\in X(k)$ can be $v$-adically close to at most $m$ of the divisors $D_{i}$. This implies that there exists a constant $c$ such that for any $P\in X(k)\setminus\cup_{i=1}^{q}\operatorname{Supp}D_{i}$ and any $v\in S$, there are indices $i_{1},\ldots,i_{m^{\prime}}\in I\subset\{1,\ldots,q\}$, $m^{\prime}\leq m$ such that
(18) $\log\prod_{i\in I}\boxdot_{D_{i},v}^{c_{i}\gamma_{D_{i}}}(P)<\log\prod_{j=1}^{m^{\prime}}\boxdot_{D_{i_{j}},v}^{c_{i_{j}}\gamma_{D_{i_{j}}}}(P)+c.$
After reindexing the divisors $D_{i,v}$, it suffices to prove a slightly weaker inequality that contains $m$-index or fewer divisors instead of $m+1$-index that in our Corollary 3.6. Note that there are only finitely many possibilities for the subset $I$ and the divisors $\left\{D_{i_{m^{\prime}}}\right\}$, then iteratively applying the Corollary 3.4 finite times, there exist a proper Zariski-closed subset $Z\subset X$ and $\epsilon>0$ such that
$\log\prod_{v\in S}\max_{I}\prod_{i\in I}\boxdot^{c_{i}\gamma_{D_{i}}}_{D_{i,v}}(P)<\left((n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}+\epsilon\right)h_{A}(P)+O(1).$
Integrated with the reformulation of Part II’s result implies our result immediately, which is a relatively weaker but more easily expressed generalization. ∎
Appendix A ascertained factor
It is notable that once we set $c_{i,v}=1$ for all $i$ and $v\in S$, we can construct the following exceptional case based on the Example 1.3 in [HL25], which demonstrates the sharpness of our results and provides a crucial illustration for a proposition:
Set $m$-divisors (independent of the $v\in S$) intersect at a point $P_{0}$ but otherwise intersect generally. Additionally, choose a family of distinct divisors $\left\{D_{i}\right\}$ ($i=1,\ldots,k\in\mathbb{Z}^{+}$), each individual intersecting the above $m$-divisors generally and mutually intersect generally. Let $D_{m+i_{s}}=D_{i}$ with $s=1,\ldots,\operatorname{codim}D_{i}$. Let these $q=m+\sum_{i=1}^{k}\operatorname{codim}D_{i}\in\mathbb{Z}^{+}$ divisors together form $\mathcal{D}_{v}$. It can plainly set $\mathcal{D}_{v}$ in $m$-subgeneral position as initial conditions. According to our setup, there exists the maximum as follows:
$\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}=\frac{m}{n}.$
Inducing the following factor formula by invoking our results in the proof section:
$a_{\delta}=\left[(n+1)\max_{\begin{subarray}{c}v\in S\\ \varnothing\subsetneq W\subsetneq X\end{subarray}}\frac{\alpha_{v}(W)}{\operatorname{codim}W}\right]_{w.r.t-\delta}=\max_{1\leq j\leq\delta n}\left(\frac{\delta m}{j}\right)(j+1)=2\delta m.$
Thus the Theorem 1.7 yields the trivial result ($\epsilon_{D_{i}}(A)\Leftrightarrow\gamma_{D_{i}}=\frac{1}{d_{i}}$):
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{q}\boxdot^{\frac{1}{d_{i}}}_{D_{i,w}}(P)<\left(2\delta m+\epsilon\right)h_{A}(P)+O(1)$
for all points $P\in X(\bar{k})\setminus\operatorname{Supp}\sum_{i=1}^{q}D_{i}$ and $\epsilon>0$, as desired.
Consider a line $L$ that passing through $P_{0}$ and is not contained within any of the divisors $D_{1},\ldots,D_{m}$. Such a line $L$ intersects $D_{1},\ldots,D_{q}$ in at most two distinct points and contains an infinite Zariski dense set of $(\sum_{i=1}^{q}D_{i},S)$-integral points. Consequently, the factor $2\delta m+\epsilon$ cannot be replaced by any smaller value.
If $m=n\geq 2$ (degenerate into general position), then we certainly have
$C(\delta,1,n,n)\leq B(1,\delta n,\delta n)\leq\max_{1\leq j\leq\delta n}A(\delta n,j)\leq\max_{1\leq j\leq\delta n}\left(\frac{\delta n}{j}\right)(j+1)=2\delta n.$
Under our assumption, we assert that
$C(\delta,1,n,n)=B(1,\delta n,\delta n)=2\delta n,$
based on the inequalities above and the following example proposed by Levin.
Example A.1.
(Example 7.4, [Lev14]) Let $\delta$ and $n\geq 2$ be positive integers. Consider $D_{1},\ldots,D_{2\delta n}\subset\mathbb{P}^{n}$ as hyperplanes in general position over a number field $k$ such that the intersection $\bigcap_{i=(s-1)n+1}^{sn}D_{i}=\{P_{s}\}$ consists of a single point for $s=1,\ldots,2\delta$, where $P_{1},\ldots,P_{2\delta}$ all lie on a line $L$ over $k$. Let $S$ be a finite set of places of $k$ containing the archimedean places, it can be verified that there exists a finite set of places $S$ of $k$ and an infinite set of points $R\subset L(\overline{k})\subset\mathbb{P}^{n}(\overline{k})$ such that
$\log\prod_{v\in S}\prod_{\begin{subarray}{c}w\in M_{k(P)}\\ w\mid v\end{subarray}}\prod_{i=1}^{2\delta n}\boxdot_{D_{i,w}}(P)=2\delta nh(P)+O(1)$
and $[k(P):k]\leq\delta$ for all $P$ in $R$.
This example implies that $C(\delta,1,n,n)\geq 2\delta n$, thus we have our conclusion.